### v1

# HR-INR: Continuous Space-Time Video Super-Resolution via Event Camera

Abstract
Continuous space-time video super-resolution (C-STVSR) aims to simultaneously enhance video resolution and frame rate at an arbitrary scale. Recently, implicit neural representation (INR) has been applied to video restoration, representing videos as implicit fields that can be decoded at an arbitrary scale. However, the highly ill-posed nature of C-STVSR limits the effectiveness of current INR-based methods: they assume linear motion between frames and use interpolation or feature warping to generate features at arbitrary spatiotemporal positions with two consecutive frames. This restrains C-STVSR from capturing rapid and nonlinear motion and long-term dependencies (involving more than two frames) in complex dynamic scenes. In this paper, we propose a novel C-STVSR framework, called HR-INR, which captures both holistic dependencies and regional motions based on INR. It is assisted by an event camera – a novel sensor renowned for its high temporal resolution and low latency. To fully utilize the rich temporal information from events, we design a feature extraction consisting of (1) a regional event feature extractor – taking events as inputs via the proposed event temporal pyramid representation to capture the regional nonlinear motion and (2) a holistic event-frame feature extractor for long-term dependence and continuity motion. We then propose a novel INR-based decoder with spatiotemporal embeddings to capture long-term dependencies with a larger temporal perception field. We validate the effectiveness and generalization of our method on four datasets (both simulated and real data), showing the superiority of our method.
1 Introduction
The real world’s visual information, e.g., edge and object motion, is continuous, spanning both time and space dimensions. However, the limited I/O bandwidth and sensor size of modern systems [13, 49, 47] confines us to record videos at low frame rates and fixed resolutions. This limitation has profound repercussions across various computer vision applications, e.g., encompassing immersive experiences in virtual reality [77, 31], traffic analysis in autonomous driving [84, 81]. To address this limitation, recent research works [10, 8] have explored restoring videos with continuous resolutions and frame rates, referred to as Continuous Space-Time Video Super-Resolution (C-STVSR).
Recently, implicit neural representation (INR) has been applied to video restoration: it represents images or videos as neural fields that can be decoded at any resolution with a pointwise MLP decoder [9, 74, 38, 4, 74, 7]. One of the seminal INR approaches is LIIF [9], which is designed for arbitrary-scale image SR. This line of research soon extended to the video domain. In this context of C-STVSR, VideoINR [10] employs a fixed STVSR model [69] that extracts features from two consecutive video frames. Then, it introduces a temporal INR to generate inverse backward warping optical flow [42] to warp features. Lastly, it employs a spatial INR, similar to LIIF, to decode the RGB frame with arbitrary resolution. Building upon this, MoTIF [8] improves VideoINR by using forward motion estimation, reducing gaps and holes in the temporal INR, which are typically caused by the randomness and discontinuities associated with backward warping [22, 46]. These methods depend solely on two successive RGB frames, rendering the task of predicting inter-frame motions ill-posed. Consequently,, it becomes challenging to accurately capture highly dynamic motion (e.g., regional high-speed or nonlinear movements) and to model long-term dependencies that extend across more than four frames.
Figure 1: With event data as guidance, our method (HR-INR) takes in videos with low frame rates and resolution (a) and produces continuous space-time videos with arbitrary frame rate and resolution (b). As shown in (c), our method is able to recover the rotation of the bicycle wheels, which is unachievable by the prior method VideoINR [10].
Motivation and Contributions. Event cameras are bio-inspired sensors, known for their high temporal resolution and low latency ($<1us$[83, 14, 63]. Recent research has demonstrated the potential of events in guiding various video super-resolution (VSR) [38, 23] and video frame interpolation (VFI) tasks [62, 61, 20, 75]. However, utilizing events to facilitate joint video super-resolution and frame interpolation is a challenging area yet to be explored.
This paper introduces HR-INR, a novel INR-based method that leverages events for jointly guiding VSR and VFI. It adeptly captures regional, rapid motion and holistic, long-range motion dependencies, as shown in Fig. 1. To capture the regional motion, we propose Temporal Pyramid Representation (TPR) to construct a time-series pyramid structure around the pivotal timestamp of events (Sec. 3.1). Different from the evenly divided timeline representations, like voxel grids [62, 61, 14], time surfaces [53], time moments [17, 38] and symmetric cumulative [57], TPR offers finer temporal granularity with less complexity and effectively captures rapid motion changes, as shown in Fig. 2 (a). Additionally, our method is capable of processing multiple frames and their associated events, which empowers it to estimate holistic, long-range motion that involves more than two frames.
Accordingly, we design two specialized feature extractors: the regional event feature extractor (RE) and the holistic event-frame feature extractor (HE), see Sec. 3.2. Both extractors are grounded in the Swin-Transformer architecture [32, 36, 33], renowned for its efficacy and efficiency in video enhancement tasks. RE is a lightweight network designed specifically for extracting local information from our event TPR. Meanwhile, HE employs a more sophisticated approach, utilizing long-term and multi-scale fusion strategies to integrate both events and frames. Consequently, our training and inference strategy requires HE to be invoked only once for multi-frame interpolation. Subsequently, the extracted features from RE and HE are fused as the output of the feature extraction module.
After fusing the regional and holistic features, we propose a novel INR-based spatial-temporal decoding module (Sec. 3.3). Our motivation is to avoid gaps and holes typically found in optical flow-based warping and multi-frame fusion, as identified in previous research [8, 10, 20, 61]. To accomplish this, we propose an implicit temporal embedding designed to transform timestamps into focused attention vectors on long-term features. This approach ensures attention is also given to long-distance dependencies, which is crucial for effectively modeling long-term temporal dependencies. Subsequently, inspired by LIIF [9], we employ spatial embedding to achieve arbitrary up-sampling in the spatial dimension.
We conducted experiments on two simulated and two real-world datasets. The results validate the superiority of our method and its excellent generalization capabilities on real-world datasets. Our approach is the first event-based method to achieve continuous space-time video super-resolution, surpassing frame-based methods, as shown in Fig. 1. It also excels in individual VSR and VFI metrics compared to previous event-based methods.
2 Related Works
Space-time Video Super-Resolution aims to enhance the resolution and frame rate of a video simultaneously [18, 27, 69, 70]. In comparison to two-stage solutions, where VFI [22, 72, 42, 43, 11] and VSR [34, 73, 76, 65, 60, 21] methods are applied sequentially, simultaneous space-time video super-resolution reduces cumulative errors and leverages the natural relations between VFI and VSR methods. Zooming Slow-Mo [69] uses temporal interpolation to generate missing frames and aligns temporal information using a deformable ConvLSTM network. Similarly, TMNet [70] extracts short-term and long-term motion cues in videos by modulating convolution kernels. However, these methods cannot simultaneously achieve spatiotemporal resolution across arbitrary scales.
INR for VFI and VSR have achieved space-time video super-resolution with arbitrary resolutions [10, 8] by learning videos implicit neural representations (INRs). These methods primarily estimate optical flows from two consecutive frames to warp features into arbitrary space-time coordinates, which are then decoded using MLP layers. However, by relying on just two consecutive frames, these methods cannot inherently model long-term motions (involving three or more frames) and fail to accurately capture local, inter-frame, non-linear motions due to missing inter-frame information.
Event-guided VFI and VSR seek to boost performance by incorporating the biologically inspired event cameras [83]. Previous works have demonstrated the potential of event-guided VFI, which mainly focus on modeling non-linear motion with events [44, 62, 68, 61, 20, 55, 54]. EFI [44] exclusively adopts the synthesis approach for intermediate frame generation. TimeLens [62] and TimeLens++ [61] employ events to model nonlinear motion correlations, integrating both synthesis and warping-centric approaches. Building on these advancements, CBM-Net [28], introduces a motion field to handle complex movements. However, while these VFI methods utilize events to capture inter-frame motion, they fail to establish long-term dependencies beyond two frames and support simultaneous VSR. The realm of event-guided video super-resolution has also been explored. E-VSR [23] highlights that high-frequency temporal information from events is beneficial to recovering high-frequency spatial information. Like our work, EG-VSR [38] employs events to comprehend INR, allowing for video ups-sampling with arbitrary scale. However, the INR of EG-VSR cannot interpolate frames. Contrasting these methods, we pioneer using events to enable concurrent VSR and VFI across arbitrary spatial-temporal scales, i.e.,, C-STVSR.
Video Long-term Dependence Modeling is a crucial aspect of VSR and VFI. For instance, BasicVSR [5] and BasicVSR++[6] enhance VSR performance by processing multi-frame inputs through an alignment module to model long-term motion correlations. Similarly, in the VFI domain, many methods [58, 41, 80] employ RNN or LSTM to model sequences of frames, capturing long-term dependencies effectively. Furthermore, Zooming Slow-Mo [69], TMNet [70], and RSST [33] leverage multi-frame inputs in the joint task of VSR and VFI, showcasing the importance of integrating multiple frames for improved modeling of video dynamics. However, current C-STVSR methods [10, 8], and event-based VFI methods [62, 20, 61, 28], primarily rely on estimating optical flow between two consecutive frames. Therefore, they are challenging to handle multi-frames as inputs, inherently undermining their capability to model long-term dependencies.
Figure 2: Overview of our framework. The inputs are multi-frame images and their corresponding events. The output is a video with enhanced frame rates and resolutions. Firstly, events proximate to a particular time point are transformed into Temporal Pyramid Representations (TPR) to capture motion at a more granular temporal level (a). Secondly, TPRs, the comprehensive set of multi-frames and events, are directed into the feature extraction part (b). Within this part, the Regional Events Feature Extractor and the Holistic Events Feature Extractor process the input separately. Lastly, the resulting features are then fused and inputted into an INR-based spatiotemporal decoding part (c). Within this part, a temporal embedding is executed to capture features at a specific timestamp $t$, followed by spatial embedding with an up-sampling factor $s$ and decoding, culminating in the generation of frames at the desired resolution.
3 Proposed Framework
Our proposed HR-INR framework is depicted in Fig. 2. The inputs of this framework are RGB frames $\boldsymbol{I}_{in}=\{I_{in}\}_{i}^{N_{in}}\in R^{N_{in}\times H\times W\times 3}$ and associated events $\boldsymbol{E}$. $H$ and $W$ denote the spatial resolution of frames and events. $3$ means three channels of RGB. $N_{in}$ denotes the input number of frames. Furthermore, the framework outputs a video with an arbitrary frame rate and spatial resolution. In particular, we consider the output video as $\boldsymbol{I}_{out}=\{I_{out}\}_{i}^{N_{out}}$, consists of $N_{out}$ frames, each with a resolution of $(s\times H)\times(s\times W)$, where $s$ represents the up-sampling scale greater than $1$. For the output $N_{out}$ frames, we denote the time corresponding to each frame as $\boldsymbol{T}=\{t\}_{i}^{N_{out}}$. For convenience, we also record the up-sampling scale $s$ and the time $\boldsymbol{T}$ as a part of inputs. Therefore, the mapping function $f_{hr}(.)$ of C-STVSR can be described by Eq. 1.
$\boldsymbol{I}_{out}=f_{hr}\left(\boldsymbol{I}_{in},\boldsymbol{E},s,\boldsymbol{T}\right)$ (1)
Our framework comprises three main components: First, Sec. 3.1 presents the event temporal pyramid representation (TPR), capturing regional dynamic motion and edges. Second, Sec. 3.2 elaborates on the feature extraction process using regional and holistic feature extractors. Third, Sec. 3.3 describes the INR-based spatiotemporal decoding.
Input Frames and Events: Our input frames, $\boldsymbol{I}_{in}$, consist of multiple frames with timestamps normalized to the $[0,1]$ interval. We consider the events, $\boldsymbol{E}$, occurring within this time range. Each event point can be represented by $(x,y,t,p)$, signifying a change in pixel intensity at coordinates $(x,y)$ at time $t$; here, $p$ is $+1$ for increased brightness and $-1$ for a decrease. Benefiting from the event camera’s high temporal resolution and low latency ($<1us$), these event points are effectively considered continuous along the timeline. For a detailed exposition of the principles of event generation, please refer to the Suppl. Mat..
3.1 Temporal Pyramid Representation
Firstly, we represent the event stream $\boldsymbol{E}$ into a frame-like form that the network can process. To capture holistic motion, we partition all events during $[0,1]$ of the timeline into $M$ equal intervals using a voxel grid [62, 14], denotes as $\boldsymbol{E}_{v}$ with dimensions $M\times H\times W$. In practice, event representation methods like voxel grid [62, 14] and its extended structure, symmetric cumulative event representation [57], achieve time granularity by uniformly dividing time into intervals with resolutions of $1/M$.
However, for C-STVSR, capturing intricate motion and edges requires a finer granularity. Merely increasing $M$ to enhance detail sharply raises computational costs; for instance, capturing $1/1000$ second intervals within a second necessitates expanding $M$ to $1000$, which is computationally impractical. To address this, we introduce Temporal Pyramid Representation (TPR), leveraging the high temporal resolution of events while reducing representation complexity.
Our idea: The core of TPR is constructing a temporal pyramid where each successive layer’s duration is $1/r,(r>1)$ of the preceding one, leading to exponentially finer time granularity with additional layers. For instance, as illustrated in Fig. 2 (a), around any given time $t$, we define a surrounding time window $\Delta t$ and select an attenuation factor $r$. At the pyramid’s $L$-th level, the events are within the time span of $[t-\Delta t/r^{L},t+\Delta t/r^{L}]$. Each layer is further segmented into $M_{p}$ intervals, represented using a voxel grid. Accordingly, for an $L$-th layer, each layer contains $M_{p}$ moments within the TPR, and its finest time granularity, denoted by $\delta_{t}$, is as delineated in the Eq. 2:
$\delta_{t}=\frac{2\times\Delta t}{M_{p}\times r^{L}}$ (2)
Therefore, for any time $t$, we construct the corresponding TPR $E_{p}$ with shape $L\times M_{p}\times H\times W$. We record the TPRs at all target timestamp as $\boldsymbol{E_{p}}=\{E_{p}\}_{i}^{N_{out}}$.
Discussion: The time granularity $\delta_{t}$ of TPR exponentially improves with the increase in layers, $L$. For an attenuation factor of $r=3$ and a goal to detect motions as brief as $1/1000$ of a second within a $1s$ window ($12\times\Delta t=$ 1), we require only $7$ TPR layers with each layer divided into $2$ intervals ($M_{p}=2$). Consequently, a TPR with dimensions $7\times 2\times H\times W$ suffices to discern motion down to $1/1000s$. Based on the above representation, we obtained $\boldsymbol{E}_{v}$, encapsulating holistic motion, and $\boldsymbol{E}_{p}$, which focuses on regional edges and motion.
3.2 Holistic-Regional Feature Extraction
This module aims to extract features from regional TPRs $\boldsymbol{E_{p}}$, and the frame $\boldsymbol{I_{in}}$ and the holistic events $\boldsymbol{E_{v}}$ for INR-based spatial-temporal decoding. Accordingly, we introduce: (1) the regional event feature extractor, $f_{re}$, for dynamic motion and edges detail capture. (2) the holistic event-frame feature extractor, $f_{he}$, for long-term motion dependencies modeling across time and space.
Figure 3: Holistic event-frame feature extractor. The down-sample module will halve the resolution. The up-sample module will double the resolution. The encoder and decoder have the same structure as Swin-Transformer [37, 35, 16].
Regional Event Feature Extraction: The input of $f_{re}$ is TPR $E_{p}\in R^{L\times M_{p}\times H\times W}$ in each timestamp $t$. Given that $f_{re}$ is invoked $N_{out}$ times, its design needs to strike a balance between efficiency and the ability to model inter-level relationships to ensure that it can capture precise regional motion and edge details effectively. Initially, our method applies a convolution layer to increase the dimensions of the feature. Then, we use four Swin Transformer Encoder Blocks (STEB) [37, 35, 16] to model the relationship between different pyramid levels. The STEB enjoys a large view field and a multi-head attention mechanism that proficiently models distant dependencies and conveys edge information across various pyramid levels, effectively capturing short-term motion. Notably, STEB is optimized with fewer parameters, boosting computational efficiency.
Holistic Event-Frame Feature Extraction: The inputs of $f_{he}$ is $\boldsymbol{I_{in}}$ and $\boldsymbol{E_{v}}$, as shown in Fig. 3. First, a convolutional layer processes both frames and events to increase dimensions. Motivated by the events feature manifest between successive frames to provide inter-frame motion information. We adjusted the first dimension of the event feature to $N_{in}-1$ to account for inter-frame motion losses. We then incorporate the STEB to facilitate interactions across varied levels and spatial domains. To minimize computational overhead and expand the receptive field, we integrated a down-sampling module between STEBs, forming a multi-scale encoder. Each down-sampling iteration halves the resolution while maintaining the channel dimensions. After three iterations, the feature resolution reduces to $1/8$ of its initial size, enlarging the receptive field. We then employ a Swin Transformer Decoder Block [37, 35, 16] to fuse features at matching resolutions. Each of the first three STEBs is followed by an upsampling process, which doubles the resolution while maintaining the channel count. Ultimately, this process outputs the feature $F_{g}$. For more details, see Suppl. Mat..
Notably, to output $N_{out}$ frames, the holistic event-frame feature extractor $f_{he}$ is called once, capturing the comprehensive feature $F^{g}$. Subsequently, for each time $t$, $f_{re}$ extracts regional features $F_{t}^{l}$ from each TPR $E_{p}\in\boldsymbol{E_{p}}$. For each regional feature $F_{t}^{l}$, we use addition and Cov1×$\times 1$ 1 operation as fusion function $f_{fu}$ to fuse with holistic feature $F^{g}$ to obtain the output $R_{t}$. For each time $t$, the whole process can be described by Eq. 3, where $E_{p}\in\boldsymbol{E_{p}}$ is the TPR at the specific timestamp $t$.
$\begin{split}R_{t}=&f_{fu}\left(F^{g},F_{t}^{l}\right);F^{g}=f_{he}\left(\boldsymbol{I_{in}},\boldsymbol{E_{v}}\right);F_{t}^{l}=f_{re}\left(E_{p}\right)\end{split}$ (3)
Figure 4: Temporal embedding. Given the input time $t\in[0,1]$, the output is the temporal attention $E_{t}$ derived from a two-layer MLP. (b) presents a visualization of the trained $E_{t}$ during $[0,1]$ on real-world dataset [61].
3.3 INR-based Spatial-Temporal Decoding
In this section, we employ INR-based spatial-temporal decoding to effectively retrieve RGB frames at any desired time and resolution. To achieve this, we leverage a temporal INR to generate features at any timestamp and a spatial INR to upscale the features to any spatial resolution.
Temporal Embedding: We utilize learned temporal embedding as attention vectors to aggregate the fused feature in the channel dimension in a time-specific manner. At a given timestamp $t$, we first use a two-layer MLP to increase its dimension to $C_{t}$, resulting in a temporal attention vector (as illustrated in Fig. 4 (a)). The visualization of the learned temporal attention is depicted in Fig. 4 (b), exhibiting variations across both time and channel dimensions. This attention vector is then multiplied directly with $R_{t}$ to generate the temporal embedded feature $R_{ts}$. This temporal INR allows for a larger temporal perception field without the need for estimating optical flows. Next, a $1\times 1$ convolution is applied to compress $R_{ts}$ to $C_{ts}$ dimensions, reducing the complexity of the spatial embedding and decoding.
Spatial Embedding and Decoding: To upscale the temporal embedded features to any desired spatial scale, we utilize a similar approach to previous works [10, 8, 9]. We query the four nearest neighbors in the temporal embedded feature for each spatial coordinate and concatenate these with their distances for spatial embedding. A four-layer MLP decoder computes the RGB values, which are then aggregated through area-weighted interpolation for arbitrary-scale super-resolution. Similar to [38], we use the Charbonnier loss [30] as the fundamental loss function for training.
Figure 5: 7-skip frame interpolation visualization results in Adobe240 dataset [56]. Our method (c) more effectively captures the rotating wheels compared to the VideoINR (b), which tends to show noticeable holes and gaps. Green circles highlight obvious holes and gaps.
4 Experiments
To facilitate a comprehensive comparison between the frame-base [10, 8] and event-guided methods [38, 62, 61, 23], we employed two simulated datasets [56, 40] and two real-world datasets [61, 51] for our experiments (1) Adobe240 dataset [56] includes 133 videos, each with $720\times 1280$ resolution and a $240fps$ frame rate. Following established protocols [10, 8, 70], the dataset is divided into 100 training, 16 validation, and 17 testing subsets. We employed the widely-used event simulation method vid2e [15], which accounts for real noise distribution, enhancing our model’s robustness and generalization capabilities. (2) GoPro dataset [40] featuring the same resolution and frame-rate with Adobe240 dataset, comprises 22 training and 11 test videos. Given its compact size, previous studies [8, 10] have primarily employed the test set for quantitative analysis. We follow this to be consistent with established practice. For datasets (1) and (2) during 7-skip frame interpolation, we input 4 frames where each pair of adjacent frames is separated by 7 frames, allowing a total of 25 frames to serve as GTs. (3) BS-ERGB [61] recorded using a spectroscope, includes real-world paired events and frames. Previous studies, e.g., TimeLens++[61], initially pre-trained on the Vime90k simulation dataset[72] before fine-tuning on BS-ERGB. In contrast, we opted to pre-train our model on the Adobe240 dataset before fine-tuning. Notably, the Vime90k dataset is larger in scale than the Adobe240 dataset. During fine-tuning, we also used perceptual loss [25] with weight $0.1$ to be consistent with previous methods [62, 61] for fair comparison. (4) CED [51] is a real-world dataset in VSR. To fairly compare the previous research [23, 38], only this data set is used for training for VSR comparison.
Implementation Details: Our model is trained using Pytorch [48], employing the Adam optimizer [29]. Referring to the VideoINR [10], our training consists of two stages. (1) Train frame interpolation under a fixed spatial up-sampling (4×4$\times$4 ×), over 70 epochs, starting with a learning rate of $45e-$ 4. (2) Train frame interpolation under random space upsampling rate in $\mathcal{U}(1,8)$, spanning 30 epochs with the learning rate $55e-$ 5. We randomly choose 20 frames from a pool of 25 frames as ground truth. Data augmentation is implemented via Random Crop, extracting $512\times 512$ areas from frames and events, with the input resolution dynamically determined by a randomly chosen upsampling ratio $s$. To optimize memory usage and accelerate speed, we implemented the mixed precision strategy [39, 12]. All experiments are performed on an NVIDIA A800 computing card. Please refer to the Suppl. Mat. for more details.
Evaluation: To ensure the fairness, we adopted PSNR [78], SSIM [67], and LPIPS [78] as quantitative evaluation metrics. Aligning with prior works [10, 8] for consistency, we use only the $Y$-channel for GoPro and Adobe datasets and all three RGB channels for BS-ERGB and CED datasets.
Table 1: Quantitative metrics (PSNR/SSIM) with 7-skip VFI and 4×4$\times$4 × VSR. Center Average remain consistent with previous work [10, 8].
VFI VSR Params ($M$) GoPro-Center GoPro-Average Adobe-Center Adobe-Average
Bicubic 19.8 27.04/0.7937 26.06/0.7720 26.09/0.7435 25.29/0.7279
Su-SloMo [22] EDVR [66] 19.8+20.7 28.24/0.8322 26.30/0.7960 27.25/0.7972 25.95/0.7682
BasicVSR [5] 19.8+6.3 28.23/0.8308 26.36/0.7977 27.28/0.7961 25.94/0.7679
[…]
QVI [71] EDVR [66] 29.2+20.7 27.43/0.8081 25.55/0.7739 26.40/0.7692 25.09/0.7406
BasicVSR [5] 29.2+6.3 27.44/0.8070 26.27/0.7955 26.43/0.7682 25.20/0.7421
[…]
DAIN [3] EDVR [66] 24.0+20.7 28.01/0.8239 26.37/0.7964 27.06/0.7895 26.01/0.7703
BasicVSR [5] 24.0+6.3 28.00/0.8227 26.46/0.7966 27.07/0.7890 26.23/0.7725
Zooming Slow Mo [69] 11.10 30.69/0.8847 -/- 30.26/0.8821 -/-
TMNet [70] 12.26 30.14/0.8692 28.83/0.8514 29.41/0.8524 28.30/0.8354
Video INR-fixed [10] 11.31 30.73/0.8850 -/- 30.21/0.8805 -/-
Video INR [10] 11.31 30.26/0.8792 29.41/0.8669 29.92/0.8746 29.27/0.8651
MoTIF [8] 12.55 31.04/0.8877 30.04/0.8773 30.63/0.8839 29.82/0.8750
HR-INR (Ours) 8.27 31.97/0.9298 32.13/0.9371 31.26/0.9246 31.11/0.9216
Table 2: More quantitative comparisons using PSNR/SSIM on the GoPro dataset.
Temporal Scale Spatial Scale Su-SloMo [22] + LIIF [9] DAIN [3] + LIIF [9] TMNet [70] Video INR [10] MoTIF [8] Ours
12$\times$ 4$\times$ 25.07/0.7491 25.14/0.7497 26.38/0.7931 27.32/0.8141 27.77/0.8230 28.87/0.8854
12$\times$ 6$\times$ 22.91/0.6783 22.92/0.6785 - 24.68/0.7358 26.78/0.7908 27.14/0.8173
16$\times$ 4$\times$ 24.42/0.7296 24.20/0.7244 24.72/0.7526 25.81/0.7739 25.98/0.7758 27.29/0.8556
16$\times$ 6$\times$ 23.28/0.6883 $2$ 23.86/0.7123 25.34/0.7527 26.09/0.7954
6$\times$ 1$\times$ - - - 32.34/0.9545 34.77/0.9696 38.53/0.9735
4.1 Comparison Experiments
Space-time Super-resolution: We conduct space-time super-resolution comparison experiments on the Adobe240 and GoPro datasets. We categorized the comparison methods into three groups. (I) Frame-based cascade methods: VFI methods, e.g., Super SloMo [22] and DAIN [3] followed by VSR methods, e.g., EDVR [66] and BasicVSR [5]. (II) Fixed STVSR methods: e.g., Zooming Slow-Mo [69] and TMNet [70]. (III) Frame-based C-STVSR methods: VideoINR [10] and MoTIF [8]. The numerical comparison is presented in Tab. 1. It is evident that C-STVSR methods consistently outperform cascade and fixed STVSR methods.Our method achieves the highest performance in both datasets with the smallest model size. In the GoPro dataset, our method improves the center frame by 0.93 $dB$ and 0.0415 SSIM, and on average by 2.09 $dB$ and 0.0598 SSIM compared to the best method, MoTIF [8]. Similarly, in the Adobe240 dataset, our method outperforms MoTIF [8] by 0.63 $dB$ and 0.0407 SSIM for the -center, and on -average by 1.29 $dB$ and 0.0466 SSIM. It is worth noting that the difference between our method and other methods is more pronounced for the -average frames than the -center frame. This observation suggests that our method demonstrates enhanced temporal stability and superior adaptability across varying interpolation intervals, an advantage not shared by previous methods [10, 8]. For more analysis, please refer to Sec. 4.2 and Fig. 8 in appendix.
Table 3: Temporal super-resolution results, i.e.,, VFI, on BS-ERGB dataset [61].
1-skip 3-skip
Methods Params ($M$) Event PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
FLAVR [26] - 25.95 - 0.086 20.90 - 0.151
DAIN [3] 24.0 25.20 - 0.067 21.40 - 0.113
Super SloMo [22] 19.8 - - - 22.48 - 0.115
QVI [71] 29.2 - - - 23.20 - 0.110
TimeLens [62] 72.2 28.36 - 0.026 27.58 - 0.031
TimeLens++ [61] 53.9 28.56 - 0.022 27.63 - 0.026
CBMNet [28] 15.4 29.32 0.815 - 28.46 0.806 -
CBMNet-Large [28] 22.2 29.43 0.816 - 28.59 0.808 -
HR-INR (Our) 8.3 29.66 0.828 0.011 28.59 0.814 0.021
Tab. 2 presents additional comparative experiments for arbitrary spatial and temporal super-resolution. Our method consistently outperforms other methods, even in extreme space-time upsample scales, such as 16×16$\times$16 × temporal upscale and 6×6$\times$6 × spatial upscale. The visualization results, Fig. 1 (c) and Fig. 5, also demonstrate that our method effectively models regional nonlinear motion, e.g., the wheel rotation — a capability not achieved by previous method VideoINR [10]. For more visualization results, please refer to the Suppl. Mat.
Figure 6: 3-skip frame interpolation visualization results in BS-ERGB dataset [61]. Our method accurately captures local non-linear motion (e.g., balls in (1.d) and (2.d)), surpassing TimeLens [61], which exhibits ghosting and holes (green circles). A yellow arrow shows TimeLens’s inaccurate ball positioning.
Separate Comparison of Event-based VFI and VSR: We also compared our method with previous approaches in separate VFI [61, 62, 28] and VSR [38, 23] tasks. The results can be seen in Tab. 3 and Tab. 4 respectively. In the VFI task, our method surpasses TimeLens++ [61] by 1.1 $dB$ for 1-skip and 0.17 $dB$ for 3-skip VFI, and CBMNet [28] by 0.23 $dB$ for 1-skip. Additionally, our model’s size is merely $1/7$ and $1/3$ that of TimeLens++ [61] and CBMNet [28], respectively. In the visualization aspect, our approach excels in modeling non-linear motion and long-term dependencies, as evident in Fig. 6 in appendix. It precisely forecasts the positions of small balls (Fig. 6 (1.d)) and the soccer ball (Fig. 6 (2.d)) at intermediate timestamps, outperforming previous methods.
For the VSR task, our method outperforms EG-VSR [38] by 1.03 $dB$ and 0.447 SSIM for 4 $\times$ super-resolution, and by 3.02 $dB$ and 0.0134 SSIM for 2 $\times$ super-resolution. Fig. 7 demonstrates our model’s effective VSR performance on the real-dataset CED, highlighting its ability to reproduce sharp edges and robustness to noise. For more visualization results, see the Suppl. Mat.
Table 4: Spatial super-resolution results on CED dataset. * denotes the values obtained from the official pre-trained models.
      4×4$\times$4 ×       2×2$\times$2 ×
      Methods       Params ($M$)       Events       PSNR $\uparrow$       SSIM $\uparrow$       PSNR $\uparrow$       SSIM $\uparrow$
      DUF [24]       1.90       ✗       24.43       0.8177       31.83       0.9183
      TDAN [60]       1.97       ✗       27.88       0.8231       33.74       0.9398
      SOF [64]       1.00       ✗       27.00       0.8050       31.84       0.9226
      RBPN [19]       12.18       ✗       29.80       0.8975       36.66       0.9754
      BasicVSR++ [6]       7.30       ✗       14.76       0.1641       -       -
      VideoINR* [10]       11.31       ✗       25.53       0.7871       26.77       0.7938
      E-VSR [23]       412.42       ✓       30.15       0.9052       37.32       0.9783
      EG-VSR [38]       2.45       ✓       31.12       0.9211       38.69       0.9771
      HR-INR (Our)       8.27       ✓       32.15       0.9658       42.01       0.9905
Figure 7: 4×4$\times$4 × video super-resolution visualization results in CED dataset [52].
Table 5: Ablation studies in Adobe-Average [56] (4×4$\times$4 × and 7-skip). The †  symbol marks the line for comparison with other lines.
Case Events TPR Temporal Embedding Temporal Dim ($C_{t}$) Input Frames PSNR $\uparrow$ SSIM $\uparrow$
Case#1 Learning 640 4 26.84 ( $-4.27$) 0.8366 ( $-0.0850$)
Case#2 Learning 640 4 29.69 ( $-1.41$) 0.8974 ( $-0.0242$)
Case#3 † Learning 640 4 31.11 0.9216
Case#4 Sinusoid 640 4 30.42 ( $-0.69$) 0.9120 ( $-0.0096$)
Case#5 Learning 320 4 28.44 ( $-2.67$) 0.8700 ( $-0.0516$)
Case#6 Learning 640 2 30.41 ( $-0.70$) 0.9151 ( $-0.0065$)
Case#7 Learning 640 3 30.72 ( $-0.39$) 0.9174 ( $-0.0042$)-
Table 6: Ablation studies for TPR levels and moments in Adobe-Average [56] (4×4$\times$4 × and 7-skip). The †  symbol marks the line for comparison with other lines.
      Case       TPR Level ($L$)       TPR Moments ($M_{p}$)       PSNR $\uparrow$       SSIM $\uparrow$
      Case#1       3       3       29.93 ( $-1.18$)       0.9011 ( $-0.0205$)
      Case#2       5       3       30.32 ( $-0.79$)       0.9165 ( $-0.0051$)
      Case#3       7       3       30.78 ( $-0.33$)       0.9187 ( $-0.0029$)
      Case#4 †       7       9       31.11       0.9216
      Case#5       7       18       31.18 ( $+0.07$)       0.9228 ( $+0.0012$)
Table 7: Ablation in BS-ERGB (TimeLens++) dataset [61].
1-skip 3-skip
TPR PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
28.25 0.8187 0.018 26.65 0.7867 0.039
29.66 ( $+1.41$) 0.8281 ( $+0.0094$) 0.011 ( $-0.007$) 28.59 ( $+1.94$) 0.8140 ( $+0.0273$) 0.021 ( $-0.018$)
4.2 Ablation and Analytical Studies
Ablation and analytical studies conducted on the Adobe240 [56] and BS-ERGB [61] datasets unveiled several critical insights. On the Adobe240 dataset, we executed simultaneous 7-skip VFI and 4×4$\times$4 × VSR tests, as shown in Tab. 5 and Tab. 6, while on the BS-ERGB dataset, 1-skip and 3-skip VFI were performed in Tab. 7. Events Gain: Tab. 5-Case#1 shows that with events input replaced as zero and unchanged network architecture, PSNR and SSIM significantly drop, highlighting the importance of events for temporal motion learning. Adding events alone substantially raised PSNR by $2.85dB$ and SSIM by $0.06$. Event TPR Enhancement: Incorporating event TPR enhances performance, with PSNR and SSIM increasing with layer count, peaking at seven layers, as shown in Tab. 5-Case#2 and Tab. 6. However, expanding ETPR moments from three to seven only improves PSNR by $0.33dB$, with minimal gains beyond that, suggesting a plateau in performance benefits. The TPR also demonstrates enhancement on the BS-ERGB dataset, as shown in Tab. 7, yielding an increase of $0.92dB$ for 1-skip and an improvement of $1.13dB$ for 3-skip. This indicates that the benefits of TPR become more pronounced with the increase in the number of skips. Time Embedding Method: Table 5-Case#3#4 shows Sinusoid embedding’s results. It’s outperformed by learning-based methods, confirmed by previous research [59, 50, 82, 1], due to their superior high-frequency information capture. Temporal Dimension Impact: The INR temporal dimension significantly influences performance. Lowering the dimensions from 640 to 320 degrades performance in Tab. 5-Case#3#5, suggesting a reduction in temporal detail capture. Conversely, expanding the dimension to 960 poses instability risks (e.g., $NAN$ errors). This highlights the need to balance dimensionality and training stability. Input Frames: Tab. 5-Case#3#6#7 illustrates the impact of varying input frame counts on the final results. We observed a performance decrease of 0.70 $dB$ and 0.39 $dB$ when inputting two and three frames, respectively, compared to four frames. This indicates a clear advantage of multi-frame inputs in modeling longer-term dependencies. Moreover, even with two frames, our method also outperforms previous works [10, 8].
5 Conclusion
Our work introduced the first event-guided continuous space-time video super-resolution method. The main contributions are: (I) Event temporal pyramid representation for capturing short-term dynamic motion; (II) A feature extraction process combining holistic and regional features to manage motion dependencies; (III) A spatiotemporal decoding process based on implicit neural representation, avoiding traditional optical flow and achieving stable frame interpolation through temporal-spatial embedding. Limitations and Future Work: Training a robust C-STVSR network requires extensive time and computational resources. Addressing these challenges remains an important area for future research. Due to space limitations, please refer more details, analysis and visualization results in the appendices.
Appendix Overview
The appendix of this paper consists of five sections:
  1. 1.
    (Sec. A) More Details about the Event Generation Model:
    This section introduces the event generation model, providing insights into why events are effective for C-STVSR.
  2. 2.
    (Sec. B) More Details about the Network Structure:
    Detailed information on the network architecture is provided, including comprehensive descriptions of the inputs, outputs, and intermediate processes of each module.
  3. 3.
    (Sec. C) More Details about the Experimental Settings:
    This section describes additional experimental settings to ensure the accurate replication of our method.
  4. 4.
    (Sec. D) Additional Experimental Results and Analysis :
    Further experimental analysis and results are offered, demonstrating the superiority of our approach.
  5. 5.
    (Sec. E) Additional Visualization Results:
    This section presents more visual materials, including images and videos.
Appendix A More Details about Event Generation Model
Events are discrete points capturing the positive and negative changes in pixel brightness. Their generation hinges on brightness alterations within the logarithmic domain. Specifically, an event point $e=(x,y,t,p)$ is triggered and logged upon meeting certain criteria. Suppose $L(x,y,t)$ represents the brightness at point $(x,y)$ at any given time $t$. The event is recorded if the absolute difference $\Delta L=log\left(L(x,y,t)\right)-L\left(x,y,t-\Delta t\right)$ surpasses a predetermined threshold $C$, as formulated as Eq. 4.
$p=\left\{\begin{aligned} +1,&\Delta L>C\\ -1,&\Delta L<-C\end{aligned}\right.$ (4)
Utilizing Eq. 4, the processing pipeline for a specified pixel at coordinates $(x,y)$ at any given time $t$ and $t^{\prime}$ can be delineated by Eq. 5.
[…]
Utilizing Eq. 5, coupled with corresponding events, the intensity frame at a given time enables the computation of intensity frames for alternate times, facilitating video frame interpolation. However, events typically contain noise, and employing Eq. 5 directly cannot produce optimal outcomes [45]. Numerous studies [44, 62, 79] have demonstrated the effectiveness of employing neural networks to guide event-based modeling in robust frame interpolation. Additionally, the high temporal resolution of events aids in high spatial resolution conversion, a finding corroborated by prior research [38, 23]. In conclusion, comprehending the event generation model reveals its substantial benefits for frame interpolation and VSR tasks, making it a natural guide for continuous space-time video super-resolution (C-STVSR) tasks.
Appendix B More Details about Network Structure
Owing to space constraints in the main text, this section provides a more detailed account of the network’s results. The description begins with the Regional Event Feature Extraction component, covering its preprocessing steps and the architecture of the Swin Transformer Encoder Blocks (STEB). Subsequently, the Holistic Event-Frame Feature Extraction is detailed, which, besides STEB, incorporates the Swin Transformer Decoder Block (STDB). In summary, this section elaborates on the preprocessing for both feature extraction components and details the structures of STEB and STDB.
B.1 Regional Event Feature Extraction:
This section delves into Regional Event Feature Extraction, denoted as $f_{re}$. $f_{re}$ receives Temporal Pyramid Representation (TPR) $E_{p}$ corresponding to timestamp $t$, formatted as $R^{L\times M_{p}\times H\times W}$. Here, $L$ signifies the number of TPR layers, while $M_{p}$ indicates the count of moments within each layer. $f_{re}$ involves two primary stages: preprocessing and feature extraction via STEB. Specifically, this preprocessing first increases its dimension to $C_{r}$ through a $1\times 1$ convolution operation. This preprocessing alters the TPR’s shape to $L\times C_{r}\times H\times W$. Feature extraction is then performed using STEB, where both the input and output retain their shape throughout the process.
Swin Transformer Encoder Blocks (STEB): The STEB structure, proven effective in video super-resolution and frame interpolation [37, 35, 16], is adopted in our design. The input and output shapes of STEB are represented as $L\times C\times H\times W$ for clarity. Initially, for a window size of $M\times M$, specifically $4\times 4$ in our implementation, the input is partitioned into disjoint windows of $(M\times M)\times N\times(H/M)\times(W/M)\times C$ dimensions. Then, each window is compressed to form a feature map of shape $(M\times M)\times(N\times H\times W/M^{2})\times C$. Following this, Layer Normalization [2] and window-based multi-head self-attention [35, 16] are computed for each window, succeeded by further transformation via another Layer Normalization and a Multi-Layer Perception. Shifted window-based multi-head self-attention [35, 16] is then employed to establish cross-window connections. After applying one STEB structure, a second STEB is introduced with an identical configuration, except the input feature window is offset by $(M/2)\times(M/2)$. In total, four STEBs are utilized for comprehensive feature extraction in the Regional Event Feature Extraction.
B.2 Holistic Event-Frame Feature Extraction:
This section details the Holistic Event-Frame Feature Extraction ($f_{he}$) module. The $f_{he}$ module processes inputs $\boldsymbol{I_{in}}$ and $\boldsymbol{E_{v}}$ to generate the output $F^{g}$. Input $\boldsymbol{I_{in}}$ has a shape of $N\times 3\times H\times W$, whereas $\boldsymbol{E_{v}}$ is structured as $M\times H\times W$, with $M$ indicating the count of moments derived from the events. The output, $F^{g}$, has a structure of $C_{g}\times H\times W$. Initially, $f_{he}$ transforms $\boldsymbol{I_{in}}$ and $\boldsymbol{E_{v}}$. $\boldsymbol{I_{in}}$ and $\boldsymbol{E_{v}}$ are transformed into features with dimensions $N\times C\times H\times W$ and $N-1\times C\times H\times W$ using $1\times 1$ convolution and reshaping, followed by processing through a STEB module. The process then employs a combination of downsampling and STEB to extract features across multiple scales, each possessing a distinct view field. Notably, STEB maintains the same resolution and channel count in both its input and output. Next, the Swin Transformer Decoder (STDB) is utilized for fusing and decoding the dual-modal features. Each STDB block receives three types of inputs: the preceding STDB’s output and the outputs from the $\boldsymbol{I_{in}}$ and $\boldsymbol{E_{v}}$ encoder sections, all at a matching resolution. In practice, the outputs from the $\boldsymbol{I_{in}}$ and $\boldsymbol{E_{v}}$ encoders are concatenated, transformed into queries via an MLP layer, and then fused utilizing a multi-head attention mechanism. Notably, the first STDB module lacks a preceding STDB output, hence defaulting this input to zero. The final output, $F^{g}$, is derived from the last STDB’s output, post reshaping and $1\times 1$ convolution. In summary, $f_{he}$ employs a sequence of intricate yet efficacious transformation and fusion procedures, designed to extract multifaceted features from inputs $\boldsymbol{I_{in}}$ and $\boldsymbol{E_{v}}$ for subsequent tasks in video super-resolution and frame interpolation.
Appendix C More Details about Experiments Setting
C.1 More Details about Datasets
To facilitate a comprehensive comparison between the frame-base [10, 8] and event-guided methods [38, 62, 61, 23], we employed four datasets: two simulated (Adobe240 [56] and GoPro [40]) and two real-world (HS-ERGB [61] and CED [51]) for our experiments.
1) Adobe240 Dataset [56]: This dataset comprises 133 videos, each with a resolution of $720\times 1280$ and a frame rate of 240. We follow [10, 8, 70] to split this dataset into 100 training, 16 validation, and 17 test sets. Upon frame extraction, we employed the widely-used event simulation method vid2e [15], which accounts for real noise distribution, ensuring robust neural network training with enhanced generalization. In generating the inputs and ground truth of the network, we adopted and extended the previous works [8, 10] to accommodate multi-frame input. Specifically, input and ground truth (GT) frames are selected via sliding windows. We define the window size as $W$, the number of input frames as $N_{in}$, and the interval as $S$. They interrelate as: $W=(N_{in}-1)*(S+1)+1$ For instance, with 4 frames input at 7-frame intervals, 25 frames are chosen. The 1st, 9th, 17th, and 25th frames become the input after down-sampling, while frames 1-25 serve as the GT. We adopted two strategies in line with prior works [10, 8]: I) A fixed magnification set at 4×4$\times$4 × the input resolution, and II) A variable enlargement strategy, wherein the scaling factor is governed by a $\mathcal{U}(1,8)$ distribution.
2) GoPro Dataset [40]: Both the GoPro and Adobe240 datasets share a resolution of $720\times 1280$ and a frame rate of 240 $fps$. However, the GoPro dataset is more compact, encompassing 22 training videos and 11 for testing. Owing to this, prior research [8, 10] predominantly utilized the GoPro test set for quantitative evaluations, neglecting its training set. In the interest of fairness, we adopted the same approach.
3) BS-ERGB [61]: This dataset, captured with a spectroscope, comprises paired events and RGB frames from real-world scenarios. It has a resolution of $970\times 625$ with RGB frames captured at 28 fps. Of the 123 videos in the dataset, 47 are allocated for training, 19 for validation, and 26 for testing. Each video contains between 100 to 600 frames. Despite being a real-world dataset, its size is inadequate for standalone training of a frame interpolation network. In previous work, e.g.TimeLens++ [61], the Vime90k [72] simulation dataset was initially utilized for pre-training, followed by fine-tuning using this real-world dataset. We, on the other hand, opted for a model pre-trained on Adobe240 for our fine-tuning.
4) CED dataset [51]: This is another real-world dataset in SR research where both frames and events exhibit a resolution of $346\times 260$. Following the previous research [23], we conducted preprocessing on this dataset.
C.2 More Details about Implementation Details
Our experiments encompass four datasets, compared against a variety of benchmark methods. To guarantee fairness in these comparisons, we provide additional details regarding our experimental approach. For the Adobe240 and GoPro datasets, our training and testing protocols adhere to methods established in prior research [8, 10]. Specifically, our approach diverges from prior studies in handling the smaller-scale BS-ERGB dataset. Unlike previous methods that pre-trained on the larger Vime90k dataset before fine-tuning on BS-ERGB, we opted for a different strategy. Our method involves pre-training on the Adobe240 dataset. During this phase, we input four video frames, with a gap of seven frames between each pair. From a total of 25 frames, 20 are randomly chosen as the ground truth. In the pre-training phase, our focus is solely on modules related to frame insertion, omitting any upsampling procedures. Consequently, the spatial embedding aspect is excluded, and the decoding is performed using only a four-layer MLP decoder. For the super-resolution experiments on the CED dataset, we deviated from pre-training and instead trained our model exclusively on the CED dataset. A more detailed comparison of the 2×2$\times$2 × VSR is shown in Tab. 8. This approach was chosen to maintain a fair comparison with other works [38, 23].
Appendix D More Experiments Results and Analysis
D.1 Stability of VFI in Various Timestamps:
Figure 8: Timestamps and corresponding PSNR for each frame during 4×4$\times$4 × VSR and 7-skip VFI on the Adobe240 dataset [56].
Figure 9: 4×4$\times$4 × VSR and 7-skip VFI visualization results. Please refer to the supplementary material for the video of this case.
Our method demonstrates temporal stability, accurately estimating motion states at each time point during frame interpolation. Specifically, whether in the GoPro-Average and Adobe-Average of Tab. 1, or the 12-skip and 16-skip tests of Tab. 2, our method significantly outperforms previous methods by at least $1.2dB$ with 4×4$\times$4 × VSR.
Fig. 8 shows the relationship between timestamps and PSNR, validating the greater stability of our method compared to VideoINR [10], especially around the $0.5s$ mark, where VideoINR experiences a notable decrease. This observation is also reflected in the visualization results of Fig. 1 and Fig. 5 in main paper and Fig. 6 and Fig. 9 in appendix. In Fig. 6 and Fig. 9, whether using real-world or simulated data, our method demonstrates superior temporal consistency, exhibiting fewer artifacts compared to previous methods such as VideoINR [10] and TimeLens [62]. This showcases the real-world effectiveness of our method’s temporal stability.
The main reason is that optical flow-based methods [10, 8, 62] suffer from instability in flow estimation at time far from the reference frames (at 00 and $1$ timestamp), which impacts the motion estimation. In contrast, this limitation does not affect our method and maintains higher stability throughout the temporal sequence.
D.2 Video Super-resolution with 2×2$\times$2 × in CED dataset
Table 8: Quantitative results (PSNR/SSIM) of the proposed our framework and other methods on the CED dataset for ×$\times 2$. Because the official training code is not available, * denoted values were acquired from the pre-trained model that the authors have released.
Clip Name TDAN [60] SOF [64] RBPN[19] VideoINR [10]* E-VSR [23] EG-VSR [38] Ours
people_dynamic_wave 35.83 / 0.9540 33.32 / 0.9360 40.07 / 0.9868 27.47 / 0.8229 41.08 / 0.9891 38.78 / 0.9794 41.50 / 0.9901
indoors_foosball_2 32.12 / 0.9339 30.86 / 0.9253 34.15 / 0.9739 26.03 / 0.7766 34.77 / 0.9775 38.68 / 0.9750 42.17 / 0.9904
simple_wires_2 31.57 / 0.9466 30.12 / 0.9326 33.83 / 0.9739 26.77 / 0.8321 34.44 / 0.9773 38.67 / 0.9815 42.21 / 0.9922
people_dynamic_dancing 35.73 / 0.9566 32.93 / 0.9388 39.56 / 0.9869 27.36 / 0.8202 40.49 / 0.9891 39.06 / 0.9798 42.02 / 0.9913
people_dynamic_jumping 35.42 / 0.9536 32.79 / 0.9347 39.44 / 0.9859 27.24 / 0.8183 40.32 / 0.9880 38.93 / 0.9792 42.09 / 0.9916
simple_fruit_fast 37.75 / 0.9440 37.22 / 0.9390 40.33 / 0.9782 27.21 / 0.8456 40.80 / 0.9801 41.96 / 0.9821 43.96 / 0.9912
outdoor_jumping_infrared_2 28.91 / 0.9062 26.67 / 0.8746 30.36 / 0.9648 26.88 / 0.8226 30.70 / 0.9698 38.03 / 0.9755 42.68 / 0.9902
simple_carpet_fast 32.54 / 0.9006 31.83 / 0.8774 34.91 / 0.9502 24.21 / 0.5909 35.16 / 0.9536 36.14 / 0.9635 39.80 / 0.9853
people_dynamic_armroll 35.55 / 0.9541 32.79 / 0.9345 40.05 / 0.9878 27.26 / 0.8193 41.00 / 0.9898 38.84 / 0.9787 41.99 / 0.9915
indoors_kitchen_2 30.67 / 0.9323 29.61 / 0.9192 31.51 / 0.9551 26.44 / 0.7502 31.79 / 0.9586 37.68 / 0.9726 41.61 / 0.9901
people_dynamic_sitting 35.09 / 0.9561 32.13 / 0.9367 39.03 / 0.9862 27.63 / 0.8230 39.97 / 0.9884 38.86 / 0.9810 41.99 / 0.9917
average PSNR/SSIM 33.74 / 0.9398 31.84 / 0.9226 36.66 / 0.9754 26.77 / 0.7938 37.32 / 0.9783 38.69 / 0.9771 42.01 / 0.9905
In Tab. 8, our proposed method demonstrates the highest performance across various clips in the CED dataset, significantly outperforming previous methods, including EG-VSR. Our approach consistently achieves higher PSNR and SSIM values, indicating superior visual quality and reconstruction accuracy. For instance, on average, our method improves PSNR by 3.32 $dB$ and SSIM by 0.0134 compared to EG-VSR. These results highlight the effectiveness of our framework in handling complex scenes, showcasing its robustness and reliability in video super-resolution tasks.
D.3 Holistic-Regional feature visualization
Fig. 10 (j, l, n) showcases TPR’s regional features, emphasizing its focus on local motion and highlighting its capability to capture local dynamics as a novel event representational form. Specifically, in a scenario where the camera is fixed while filming a basketball player, the background information is displayed in the Holistic Feature, whereas the motion information (such as the moving basketball) is emphasized in the Regional Feature. This indicates that our method can effectively capture different types of information through both Holistic and Regional features. For more cases, please refer to the example figure below and the videos in the supplementary materials.
Figure 10: Feature visualization on real data [61]: (h) shows the holistic feature, $F^{g}$, derived from multiple frames and events; (j, l, n) depict the regional features ($F_{t}^{g}$), highlighting the capability to capture local motion.
D.4 Time embedding feature visualization
Figure 11: Visualization of Temporal Embedding. Figure (a) shows the original visualization of Temporal Embedding, while figure (b) displays the results after sorting. The sorting is based on the size of the first element, arranged in ascending order. We use MLP decoding, where the order is not crucial. However, to more clearly demonstrate the outcomes of Temporal Embedding learning, we have chosen to present the sorted results.
Fig.11 illustrates the visualization of time embedding features. Compared to traditional sine-cosine embedding features, the learning-based approach performs better, as shown in Tab. 5. The visualized learning-based embedding features not only demonstrate the capability to learn periodic positional representations but also provide a richer expression of exposure information.
D.5 Bad case analysis
We have observed that our method has certain limitations in some cases (Fig. 9). For example, when restoring color information, although our model can accurately reconstruct the contours of objects, the color information is often distorted or missing. This issue primarily arises due to the lack of color information in the event stream. We believe that with future advancements in color event technology, this problem will be effectively addressed.
D.6 Inference time analysis
In Fig. 11, we analyze the inference times of three different methods. Both our method and VideoINR achieve an average frame time of less than 100 $ms$ for 34×34$\times$34 × frame interpolation. In contrast, TimeLens [62] has an average frame time of 187 $ms$, which is more than double that of our method. The tests were conducted on a high-performance computer, and each method was tested 30 times, with the final inference time being the average of these 30 trials.
Figure 12: Comparison of total and average time for 34×34$\times$34 × frame interpolation by different methods. Our method takes less time than TimeLens [62], but slightly more time than VideoINR [10].
Appendix E More Visualization Results
Additional videos have been included in the supplementary materials to provide a more comprehensive demonstration of our method’s visual results. Below, we enumerate these videos and briefly describe their key features. We then present more visualizations to demonstrate the generalization of our method on real data.
  • 1-Adobe240: This video contains the following five clips.
    • IMG_0013-7skip4xsr-Cyclist: In this video, the camera and the people in the background are in motion, creating a complex scene. Our method successfully recovers the locally moving bicycle, demonstrating exceptional video frame interpolation and super-resolution capabilities.
    • IMG_0037-7skip4xsr-
      TrafficIntersectionManyCars:
      The video demonstrates a camera with slight movement, capturing a busy intersection bustling with vehicles. Our method is capable of accurately recovering vehicles in motion within the scene, including the intricate details of rotating tires.
    • IMG_0037a-7skip-
      MovingForegroundAndBackground:
      The video includes both distant and close-up elements. In the close-up scenes, the comparative methods resulted in significant deformations and distortions.
    • IMG_0045-7skip-
      PortraitSculpture:
      This video demonstrates the effects under significant camera movement. When the camera moves rapidly, frame-based methods tend to underperform.
    • IMG_0175-7skip4xsr-
      LawnAndCar:
      The same scene occurs when the camera moves violently.
    • IMG_0175-7skip4xsr-
      TreeComplexTexture:
      This video captures leaves, demonstrating that methods based on optical flow tend to fail in the presence of complex textures.
  • 2-TimeLensPP-Ours-1: This video shows the performance of our method on real-world data sets and the visualization of features. Demonstrates that we effectively capture local motion.
  • 3-Our-vs-Timelens: This video shows the results of comparing our method with Timelens [62].
Figure 13: More visualization results on real-world data set [62].
Figure 14: More visualization results on real-world data set [62].
Figure 15: More visualization results on real-world data set [62].
Figure 16: More visualization results on real-world data set [62].
Figure 17: More visualization results on real-world data set [62].
Figure 18: More visualization results on real-world data set [62].
Figure 19: More visualization results on real-world data set [62].
Figure 20: More visualization results on real-world data set [62].

### v2

@@
-HR-INR: Continuous Space-Time Video Super-Resolution via Event Camera
+
 Abstract
@@
 Recently, implicit neural representation (INR) has been applied to video restoration, representing videos as implicit fields that can be decoded at an arbitrary scale.
-However, the highly ill-posed nature of C-STVSR limits the effectiveness of current INR-based methods: they assume linear motion between frames and use interpolation or feature warping to generate features at arbitrary spatiotemporal positions with two consecutive frames.
-This restrains C-STVSR from capturing rapid and nonlinear motion and long-term dependencies (involving more than two frames) in complex dynamic scenes.
-In this paper, we propose a novel C-STVSR framework, called HR-INR, which captures both holistic dependencies and regional motions based on INR.
+However, existing INR-based C-STVSR methods typically rely on only two frames as input, leading to insufficient inter-frame motion information.
+Consequently, they struggle to capture fast, complex motion and long-term dependencies (spanning more than three frames), hindering their performance in dynamic scenes.
+In this paper, we propose a novel C-STVSR framework, named HR-INR, which captures both holistic dependencies and regional motions based on INR.
 It is assisted by an event camera – a novel sensor renowned for its high temporal resolution and low latency.
@@
 We validate the effectiveness and generalization of our method on four datasets (both simulated and real data), showing the superiority of our method.
+The project page is available at https://github.com/yunfanLu/HR-INR.
+journal: International Journal of Computer Vision
 1 Introduction
+Figure 1: With event data as guidance, our method (HR-INR) takes in videos with low frame rates and resolution (a) and produces continuous space-time videos with arbitrary frame rate and resolution (b).
+Demonstrating effective modeling of local nonlinear motion, our method uniquely showcases this with the bicycle example in (c), a feat unachievable by the prior method - VideoINR (Chen et al, 2022).
+As shown in (c), our method is able to recover the rotation of the bicycle wheels, which is unachievable by the prior method VideoINR (Chen et al, 2022).
 The real world’s visual information, e.g., edge and object motion, is continuous, spanning both time and space dimensions.
-However, the limited I/O bandwidth and sensor size of modern systems [13, 49, 47] confines us to record videos at low frame rates and fixed resolutions.
-This limitation has profound repercussions across various computer vision applications, e.g., encompassing immersive experiences in virtual reality [77, 31], traffic analysis in autonomous driving [84, 81].
-To address this limitation, recent research works [10, 8] have explored restoring videos with continuous resolutions and frame rates, referred to as Continuous Space-Time Video Super-Resolution (C-STVSR).
-Recently, implicit neural representation (INR) has been applied to video restoration: it represents images or videos as neural fields that can be decoded at any resolution with a pointwise MLP decoder [9, 74, 38, 4, 74, 7].
-One of the seminal INR approaches is LIIF [9], which is designed for arbitrary-scale image SR.
+However, the limited I/O bandwidth and sensor size of modern systems (Delbracio et al, 2021; Parker, 2010) confines us to record videos at low frame rates and fixed resolutions.
+This limitation has profound repercussions across various computer vision applications, e.g., encompassing immersive experiences in virtual reality (Zhang, 2020; Lee et al, 2020), traffic analysis in autonomous driving (Zou et al, 2023; Zhao et al, 2019).
+To address this limitation, recent research works, e.g., VideoINR (Chen et al, 2022, 2023b), have explored restoring videos by implicit neural representation (INR), with continuous resolutions and frame rates, referred to as Continuous Space-Time Video Super-Resolution (C-STVSR).
+In practice, INR-based methods represent images or videos as neural fields that can be decoded at any resolution using a pointwise MLP decoder (Cao et al, 2023; Chen et al, 2023a).
+Once an image or video is encoded into this neural field, it effectively becomes a continuous function over space and time.
+Consequently, querying any previously unseen coordinate $(x,y,t)$ yields the corresponding RGB value, thus enabling arbitrary-scale spatial upsampling and arbitrary-frame-rate temporal interpolation.
+One of the seminal INR approaches is LIIF (Chen et al, 2021), which is designed for arbitrary-scale image SR.
 This line of research soon extended to the video domain.
-In this context of C-STVSR, VideoINR [10] employs a fixed STVSR model [69] that extracts features from two consecutive video frames.
-Then, it introduces a temporal INR to generate inverse backward warping optical flow [42] to warp features.
+In this context of C-STVSR, VideoINR (Chen et al, 2022) employs a fixed STVSR model (Xiang et al, 2020) that extracts features from two consecutive video frames.
+Then, it introduces a temporal INR to generate inverse backward warping optical flow (Niklaus and Liu, 2020) to warp features.
 Lastly, it employs a spatial INR, similar to LIIF, to decode the RGB frame with arbitrary resolution.
-Building upon this, MoTIF [8] improves VideoINR by using forward motion estimation, reducing gaps and holes in the temporal INR, which are typically caused by the randomness and discontinuities associated with backward warping [22, 46].
-These methods depend solely on two successive RGB frames, rendering the task of predicting inter-frame motions ill-posed.
-Consequently,, it becomes challenging to accurately capture highly dynamic motion (e.g., regional high-speed or nonlinear movements) and to model long-term dependencies that extend across more than four frames.
-Figure 1: With event data as guidance, our method (HR-INR) takes in videos with low frame rates and resolution (a) and produces continuous space-time videos with arbitrary frame rate and resolution (b).
-As shown in (c), our method is able to recover the rotation of the bicycle wheels, which is unachievable by the prior method VideoINR [10].
-Motivation and Contributions.
-Event cameras are bio-inspired sensors, known for their high temporal resolution and low latency ($<1us$[83, 14, 63].
-Recent research has demonstrated the potential of events in guiding various video super-resolution (VSR) [38, 23] and video frame interpolation (VFI) tasks [62, 61, 20, 75].
-However, utilizing events to facilitate joint video super-resolution and frame interpolation is a challenging area yet to be explored.
-This paper introduces HR-INR, a novel INR-based method that leverages events for jointly guiding VSR and VFI.
-It adeptly captures regional, rapid motion and holistic, long-range motion dependencies, as shown in Fig.
+Building upon this, MoTIF (Chen et al, 2023b) improves VideoINR by using forward motion estimation, reducing gaps and holes in the temporal INR, which are typically caused by the randomness and discontinuities associated with backward warping (Park et al, 2021).
+However, these methods depend solely on two successive RGB frames, rendering the task of predicting inter-frame motions ill-posed, as shown in Fig.
+1 (c).
+Consequently, it becomes challenging to accurately capture highly dynamic motion (e.g., regional high-speed or complex nonlinear movements) and to model long-term dependencies that extend across more than four frames.
+Motivation and Contributions: Event cameras are bioinspired sensors, known for their high temporal resolution and low latency (<1us<1us) (Gallego et al, 2020; Zheng et al, 2023; Lu et al, 2025b).
+Recent research has demonstrated the potential of events in guiding various video super-resolution (VSR) (Lu et al, 2023; Jing et al, 2021) and video frame interpolation (VFI) tasks (Tulyakov et al, 2021, 2022; He et al, 2022; Chen et al, 2025; Wang et al, 2025; Yan et al, 2025; Wei et al, 2025).
+Specifically, Jing et al (2021) demonstrated that the high temporal resolution of events can be converted into high spatial resolution of frames.
+Furthermore, Lu et al (2023) then extended this advantage to arbitrary-scale spatial upsampling.
+In the VFI, event streams provide motion information between frames.
+Tulyakov et al (2021, 2022) proposed dual-stream methods to leverage the temporal information from events for interpolation at any given moment.
+Kim et al (2023) modeled finer motions through motion field estimation.
+These methods have achieved significant progress in VFI and VSR tasks, and demonstrated how events can enhance both temporal information and spatial information.
+However, utilizing events to facilitate joint video super-resolution and frame interpolation remains a challenging area that has not been thoroughly explored.
+To end this, this paper introduces HR-INR, a novel INR-based method that leverages events for jointly guiding VSR and VFI through holistic and regional motion modeling.
+It adeptly achieves C-STVSR by capturing regional, rapid motion and holistic, long-range motion dependencies, as shown in Fig.
 1.
@@
 3.1).
-Different from the evenly divided timeline representations, like voxel grids [62, 61, 14], time surfaces [53], time moments [17, 38] and symmetric cumulative [57], TPR offers finer temporal granularity with less complexity and effectively captures rapid motion changes, as shown in Fig.
+Different from the evenly divided timeline representations, like voxel grids (Tulyakov et al, 2021, 2022; Gallego et al, 2020), time surfaces (Sironi et al, 2018), time moments (Han et al, 2021; Lu et al, 2023) and symmetric cumulative (Sun et al, 2022), TPR offers finer temporal granularity with less complexity and effectively captures rapid motion changes, as shown in Fig.
 2 (a).
-Additionally, our method is capable of processing multiple frames and their associated events, which empowers it to estimate holistic, long-range motion that involves more than two frames.
+Additionally, to estimate holistic, long-range motion (involving more than two frames), our model can employ more frames (e.g., four frames) and their corresponding events.
+This makes it possible to capture the extended-duration motion and improves the accuracy and continuity of motion modeling.
 Accordingly, we design two specialized feature extractors: the regional event feature extractor (RE) and the holistic event-frame feature extractor (HE), see Sec.
 3.2.
-Both extractors are grounded in the Swin-Transformer architecture [32, 36, 33], renowned for its efficacy and efficiency in video enhancement tasks.
-RE is a lightweight network designed specifically for extracting local information from our event TPR.
-Meanwhile, HE employs a more sophisticated approach, utilizing long-term and multi-scale fusion strategies to integrate both events and frames.
-Consequently, our training and inference strategy requires HE to be invoked only once for multi-frame interpolation.
+Both extractors are grounded in the Swin-Transformer architecture (Liang et al, 2021; Liu et al, 2022a), renowned for its efficacy and efficiency in video enhancement tasks.
+The regional event feature extractor is a lightweight network designed specifically for extracting local information from our event TPR.
+Meanwhile, holistic event-frame feature extractor employs a more sophisticated approach, utilizing long-term and multi-scale fusion strategies to integrate both events and frames.
+Consequently, our training and testing strategy requires HE to be invoked only once for multi-frame interpolation.
 Subsequently, the extracted features from RE and HE are fused as the output of the feature extraction module.
@@
 3.3).
-Our motivation is to avoid gaps and holes typically found in optical flow-based warping and multi-frame fusion, as identified in previous research [8, 10, 20, 61].
+Our motivation is to reduce gaps and holes typically found in optical flow-based warping and multi-frame fusion, as show in Fig.
+1 (c), which are also identified in previous research (Chen et al, 2023b, 2022; He et al, 2022; Tulyakov et al, 2022).
 To accomplish this, we propose an implicit temporal embedding designed to transform timestamps into focused attention vectors on long-term features.
-This approach ensures attention is also given to long-distance dependencies, which is crucial for effectively modeling long-term temporal dependencies.
-Subsequently, inspired by LIIF [9], we employ spatial embedding to achieve arbitrary up-sampling in the spatial dimension.
-We conducted experiments on two simulated and two real-world datasets.
+This approach is crucial for effectively modeling long-term temporal dependencies and capturing regional dynamic motion and edges, ensuring attention is also given to long-distance dependencies.
+Subsequently, we employ spatial embedding (Chen et al, 2021) to achieve arbitrary up-sampling in the spatial dimension.
+We conducted experiments on two simulated (Adobe240 (Su et al, 2017) and GoPro (Nah et al, 2017)) and two real-world datasets (BS-ERGB (Tulyakov et al, 2022) and CED (Scheerlinck et al, 2019a)).
 The results validate the superiority of our method and its excellent generalization capabilities on real-world datasets.
-Our approach is the first event-based method to achieve continuous space-time video super-resolution, surpassing frame-based methods, as shown in Fig.
+Our approach is the first event-based method to achieve continuous space-time video super resolution, surpassing frame-based methods, as shown in Fig.
 1.
@@
 2 Related Works
-Space-time Video Super-Resolution aims to enhance the resolution and frame rate of a video simultaneously [18, 27, 69, 70].
-In comparison to two-stage solutions, where VFI [22, 72, 42, 43, 11] and VSR [34, 73, 76, 65, 60, 21] methods are applied sequentially, simultaneous space-time video super-resolution reduces cumulative errors and leverages the natural relations between VFI and VSR methods.
-Zooming Slow-Mo [69] uses temporal interpolation to generate missing frames and aligns temporal information using a deformable ConvLSTM network.
-Similarly, TMNet [70] extracts short-term and long-term motion cues in videos by modulating convolution kernels.
+Since our research focuses on leveraging event data to guide the C-STVSR task, we categorize related works into five areas: (1) Frame-based VFI and VSR; (2) Frame-based Space-Time Video Super-Resolution; (3) INR for VFI and VSR; (4) Event-guided VFI and VSR; (5) Long-term Video Dependency Modeling.
+Sections (1) and (2) discuss the progress and limitations of frame-based methods, (3) highlights the role of INR in arbitrary scales, (4) covers event applications in VFI and VSR, and (5) emphasizes the importance of modeling long-term dependencies in videos.
+(1) Frame-based VFI and VSR VFI generates intermediate frames, while VSR improves video spatial resolution.
+Early frame-based works addressed these tasks separately.
+VSR focuses on modeling motion relations between multiple frames for feature alignment and resolution enhancement.
+Early methods Bao et al (2019b); Wang et al (2019) used motion compensation and deformable convolutions, while later works (Chan et al, 2021, 2022a) explored longer temporal modeling.
+VFI, on the other hand, models inter-frame motion to generate intermediate frames, with early approaches relying on warping, kernels, or optical flow (Meyer et al, 2018; Xu et al, 2019; Dutta et al, 2021; Zhang et al, 2022).
+Despite attempts to model nonlinear motion, these methods remain limited by the absence of inter-frame data.
+Recent work, such as RSST (Liang et al, 2022), improves VFI by using multi-frame inputs.
+In summary, both VSR and VFI rely on inter-frame motion modeling but are constrained by missing inter-frame motion information, making it challenging to capture complex nonlinear motion.
+(2) Frame-based Space-time Video Super-Resolution aims to enhance the resolution and frame rate of a video simultaneously (Haris et al, 2020; Kim et al, 2020; Xiang et al, 2020; Xu et al, 2021).
+In comparison to two-stage solutions, where VFI  (Jiang et al, 2018; Xue et al, 2019; Niklaus and Liu, 2020; Niklaus et al, 2017; Cheng and Chen, 2020) and VSR  (Liu et al, 2018; Yang et al, 2021; Yue et al, 2022; Wang et al, 2021; Tian et al, 2020; Isobe et al, 2020) methods are applied sequentially, simultaneous space-time video super-resolution reduces cumulative errors and leverages the natural relations between VFI and VSR methods.
+Zooming Slow-Mo (Xiang et al, 2020) uses temporal interpolation to generate missing frames and aligns temporal information using a deformable ConvLSTM network.
+Similarly, TMNet (Xu et al, 2021) extracts short-term and long-term motion cues in videos by modulating convolution kernels.
 However, these methods cannot simultaneously achieve spatiotemporal resolution across arbitrary scales.
-INR for VFI and VSR have achieved space-time video super-resolution with arbitrary resolutions [10, 8] by learning videos implicit neural representations (INRs).
-These methods primarily estimate optical flows from two consecutive frames to warp features into arbitrary space-time coordinates, which are then decoded using MLP layers.
-However, by relying on just two consecutive frames, these methods cannot inherently model long-term motions (involving three or more frames) and fail to accurately capture local, inter-frame, non-linear motions due to missing inter-frame information.
-Event-guided VFI and VSR seek to boost performance by incorporating the biologically inspired event cameras [83].
-Previous works have demonstrated the potential of event-guided VFI, which mainly focus on modeling non-linear motion with events [44, 62, 68, 61, 20, 55, 54].
-EFI [44] exclusively adopts the synthesis approach for intermediate frame generation.
-TimeLens [62] and TimeLens++ [61] employ events to model nonlinear motion correlations, integrating both synthesis and warping-centric approaches.
-Building on these advancements, CBM-Net [28], introduces a motion field to handle complex movements.
-However, while these VFI methods utilize events to capture inter-frame motion, they fail to establish long-term dependencies beyond two frames and support simultaneous VSR.
-The realm of event-guided video super-resolution has also been explored.
-E-VSR [23] highlights that high-frequency temporal information from events is beneficial to recovering high-frequency spatial information.
-Like our work, EG-VSR [38] employs events to comprehend INR, allowing for video ups-sampling with arbitrary scale.
-However, the INR of EG-VSR cannot interpolate frames.
-Contrasting these methods, we pioneer using events to enable concurrent VSR and VFI across arbitrary spatial-temporal scales, i.e.,, C-STVSR.
-Video Long-term Dependence Modeling is a crucial aspect of VSR and VFI.
-For instance, BasicVSR [5] and BasicVSR++[6] enhance VSR performance by processing multi-frame inputs through an alignment module to model long-term motion correlations.
-Similarly, in the VFI domain, many methods [58, 41, 80] employ RNN or LSTM to model sequences of frames, capturing long-term dependencies effectively.
-Furthermore, Zooming Slow-Mo [69], TMNet [70], and RSST [33] leverage multi-frame inputs in the joint task of VSR and VFI, showcasing the importance of integrating multiple frames for improved modeling of video dynamics.
-However, current C-STVSR methods [10, 8], and event-based VFI methods [62, 20, 61, 28], primarily rely on estimating optical flow between two consecutive frames.
-Therefore, they are challenging to handle multi-frames as inputs, inherently undermining their capability to model long-term dependencies.
+i.e.,C-STVSR.
 Figure 2: Overview of our framework.
@@
 Within this part, a temporal embedding is executed to capture features at a specific timestamp $t$, followed by spatial embedding with an up-sampling factor $s$ and decoding, culminating in the generation of frames at the desired resolution.
+(3) INR for VFI and VSR have achieved space-time video super-resolution with arbitrary resolutions (Chen et al, 2022, 2023b) by learning videos implicit neural representations (INRs).
+These methods primarily estimate optical flows from two consecutive frames to warp features into arbitrary space-time coordinates, which are then decoded using MLP layers.
+However, by relying on just two consecutive frames, these methods cannot inherently model long-term motions (involving three or more frames) and fail to accurately capture local, inter-frame, non-linear motions due to missing inter-frame information.
+(4) Event-guided VFI and VSR seek to boost performance by incorporating the biologically inspired event cameras (Zheng et al, 2023; Lu et al, 2025a).
+Previous works have demonstrated the potential of event-guided VFI, which mainly focus on modeling non-linear motion with events  (Tulyakov et al, 2021, 2022; He et al, 2022; Song et al, 2023; Chen et al, 2025; Wang et al, 2025; Yan et al, 2025; Wei et al, 2025).
+EFI (Paikin et al, 2021) exclusively adopts the synthesis approach for intermediate frame generation.
+TimeLens (Tulyakov et al, 2021) and TimeLens++ (Tulyakov et al, 2022) employ events to model nonlinear motion correlations, integrating both synthesis and warping-centric approaches.
+Building on these advancements, CBM-Net (Kim et al, 2023), introduces a motion field to handle complex movements.
+E-CIR (Song et al, 2022) and DeblurSR (Song et al, 2023) use events to learn a function to map time to gray values, which can speed up the prediction time of each frame.
+However, while these VFI methods utilize events to capture inter-frame motion, they fail to establish long-term dependencies beyond two frames and support simultaneous VSR.
+The realm of event-guided video super-resolution has also been explored.
+E-VSR (Jing et al, 2021) highlights that high-frequency temporal information from events is beneficial to recovering high-frequency spatial information.
+EG-VSR (Lu et al, 2023) employs events to comprehend INR, allowing for video ups-sampling with arbitrary scale.
+Contrasting these methods, we pioneer using events to enable concurrent VSR and VFI across arbitrary spatial-temporal scales, i.e.,C-STVSR.
+(5) Video Long-term Dependence Modeling is a crucial aspect of VSR and VFI.
+For instance, BasicVSR (Chan et al, 2021) and BasicVSR++(Chan et al, 2022b) enhance VSR performance by processing multi-frame inputs through an alignment module to model long-term motion correlations.
+Similarly, in the VFI domain, many methods (Suzuki and Ikehara, 2020; Nah et al, 2019; Zhang et al, 2020) employ RNN or LSTM to model sequences of frames, capturing long-term dependencies effectively.
+Furthermore, Zooming Slow-Mo (Xiang et al, 2020), TMNet (Xu et al, 2021), and RSST (Liang et al, 2022) leverage multi-frame inputs in the joint task of VSR and VFI, showcasing the importance of integrating multiple frames for improved modeling of video dynamics.
+However, current C-STVSR methods (Chen et al, 2022, 2023b), and event-based VFI methods (Tulyakov et al, 2021; He et al, 2022; Tulyakov et al, 2022; Kim et al, 2023), primarily rely on estimating optical flow between two consecutive frames, which is then used to warp these two frames to generate the intermediate frames.
+Therefore, they are challenging to handle multi-frames as inputs, inherently undermining their capability to model long-term dependencies.
 3 Proposed Framework
-Our proposed HR-INR framework is depicted in Fig.
+Preliminary: Events are discrete points that capture the positive and negative changes in pixel brightness.
+Their generation depends on brightness alterations within the logarithmic domain.
+Specifically, an event point $e=(x,y,t,p)$ is triggered and logged upon meeting certain criteria.
+Suppose $L(x,y,t)$ represents the brightness at point $(x,y)$ at any given time $t$.
+The event is recorded if the absolute difference $\Delta L=log\left(L(x,y,t)\right)-L\left(x,y,t-\Delta t\right)$ surpasses a predetermined threshold $C$, as formulated as Eq.
+1.
+$p=\left\{\begin{aligned} +1,&\Delta L>C\\ -1,&\Delta L<-C\end{aligned}\right$. (1)
+Utilizing Eq. 1, the processing pipeline for a specified pixel at coordinates $(x,y)$ at any given time $t$ and $t^{\prime}$ can be delineated by Eq. 2.
+$L(x,y,t)=I(x,y,t^{\prime})\times\exp({C\int_{t^{\prime}}^{t}p~dt})$ (2)
+Utilizing Eq.
+2 along with corresponding events, intensity frames at a given time can be used to compute frames at alternate times, facilitating video frame interpolation.
+However, events are often noisy, and directly applying Eq.
+2 may not yield optimal results (Pan et al, 2019).
+Therefore, employing neural networks for event-based frame interpolation has garnered significant attention (Paikin et al, 2021; Tulyakov et al, 2021; Zhang and Yu, 2022; Kim et al, 2023).
+Additionally, the high temporal resolution of events helps in converting to high spatial resolution, as demonstrated by previous studies (Lu et al, 2023; Jing et al, 2021).
+In conclusion, event signals play a crucial role in frame interpolation and video super-resolution tasks, making them a natural guide for continuous space-time video super-resolution.
+Overview of HR-INR: Our proposed HR-INR framework is depicted in Fig.
 2.
-The inputs of this framework are RGB frames $\boldsymbol{I}_{in}=\{I_{in}\}_{i}^{N_{in}}\in R^{N_{in}\times H\times W\times 3}$ and associated events $\boldsymbol{E}$.
+The inputs of this framework are RGB frames $\bm{I}_{in}=\{I_{in}\}_{i}^{N_{in}}\in R^{N_{in}\times H\times W\times 3}$ and associated events $\bm{E}$.
 $H$ and $W$ denote the spatial resolution of frames and events.
@@
 Furthermore, the framework outputs a video with an arbitrary frame rate and spatial resolution.
-In particular, we consider the output video as $\boldsymbol{I}_{out}=\{I_{out}\}_{i}^{N_{out}}$, consists of $N_{out}$ frames, each with a resolution of $(s\times H)\times(s\times W)$, where $s$ represents the up-sampling scale greater than $1$.
-For the output $N_{out}$ frames, we denote the time corresponding to each frame as $\boldsymbol{T}=\{t\}_{i}^{N_{out}}$.
-For convenience, we also record the up-sampling scale $s$ and the time $\boldsymbol{T}$ as a part of inputs.
+In particular, we consider the output video as $\bm{I}_{out}=\{I_{out}\}_{i}^{N_{out}}$, consists of $N_{out}$ frames, each with a resolution of $(s\times H)\times(s\times W)$, where $s$ represents the up-sampling scale greater than $1$.
+For the output $N_{out}$ frames, we denote the time corresponding to each frame as $\bm{T}=\{t\}_{i}^{N_{out}}$.
+For convenience, we also record the up-sampling scale $s$ and the time $\bm{T}$ as a part of inputs.
 Therefore, the mapping function $f_{hr}(.)$ of C-STVSR can be described by Eq.
-1.
-$\boldsymbol{I}_{out}=f_{hr}\left(\boldsymbol{I}_{in},\boldsymbol{E},s,\boldsymbol{T}\right)$ (1)
+3.
+$\bm{I}_{out}=f_{hr}\left(\bm{I}_{in},\bm{E},s,\bm{T}\right)$ (3)
 Our framework comprises three main components: First, Sec.
@@
 3.3 describes the INR-based spatiotemporal decoding.
-Input Frames and Events: Our input frames, $\boldsymbol{I}_{in}$, consist of multiple frames with timestamps normalized to the $[0,1]$ interval.
-We consider the events, $\boldsymbol{E}$, occurring within this time range.
-Each event point can be represented by $(x,y,t,p)$, signifying a change in pixel intensity at coordinates $(x,y)$ at time $t$; here, $p$ is $+1$ for increased brightness and $-1$ for a decrease.
-Benefiting from the event camera’s high temporal resolution and low latency ($<1us$), these event points are effectively considered continuous along the timeline.
-For a detailed exposition of the principles of event generation, please refer to the Suppl.
-Mat..
+Figure 3: Visualization of Event TPR across different time resolutions.
+The TPR is divided into 7 layers ($L=7$), with each layer having a time resolution that is $1/3$ of the previous layer ($r=3$).
+The time resolution for the first layer ($L_{0}$) is approximately $2\times\Delta t=1/9s$, and the time resolution for the seventh layer ($L_{6}$) is approximately $1/6561s$.
+Each layer visualizes the corresponding event data at a specific time resolution, demonstrating the effect of varying temporal resolution on event-based data.
 3.1 Temporal Pyramid Representation
-Firstly, we represent the event stream $\boldsymbol{E}$ into a frame-like form that the network can process.
-To capture holistic motion, we partition all events during $[0,1]$ of the timeline into $M$ equal intervals using a voxel grid [62, 14], denotes as $\boldsymbol{E}_{v}$ with dimensions $M\times H\times W$.
-In practice, event representation methods like voxel grid [62, 14] and its extended structure, symmetric cumulative event representation [57], achieve time granularity by uniformly dividing time into intervals with resolutions of $1/M$.
+Firstly, we represent the event stream $\bm{E}$ into a frame-like form that the network can process.
+To capture holistic motion, we partition all events during $[0,1]$ of the timeline into $M$ equal intervals using a voxel grid (Tulyakov et al, 2021; Gallego et al, 2020), denotes as $\bm{E}_{v}$ with dimensions $M\times H\times W$.
+In practice, event representation methods like voxel grid (Tulyakov et al, 2021; Gallego et al, 2020) and its extended structure, symmetric cumulative event representation (Sun et al, 2022), achieve time granularity by uniformly dividing time into intervals with resolutions of $1/M$.
 However, for C-STVSR, capturing intricate motion and edges requires a finer granularity.
@@
 2 (a), around any given time $t$, we define a surrounding time window $\Delta t$ and select an attenuation factor $r$.
-At the pyramid’s $L$-th level, the events are within the time span of $[t-\Delta t/r^{L},t+\Delta t/r^{L}]$.
+At the pyramid’s $L$-th level, the events are within the time span of $t-\Delta t/r^{L},t+\Delta t/r^{L}]$.
 Each layer is further segmented into $M_{p}$ intervals, represented using a voxel grid.
 Accordingly, for an $L$-th layer, each layer contains $M_{p}$ moments within the TPR, and its finest time granularity, denoted by $\delta_{t}$, is as delineated in the Eq.
-2:
+4.
+Therefore, for any time $t$, we construct the corresponding TPR $E_{p}$ with shape $L\times M_{p}\times H\times W$.
+A simplified example is shown in Fig.
+3.
+We record the TPRs at all target timestamp as $\bm{E_{p}}=\{E_{p}\}_{i}^{N_{out}}$.
 $\delta_{t}=\frac{2\times\Delta t}{M_{p}\times r^{L}}$ (4)
-Therefore, for any time $t$, we construct the corresponding TPR $E_{p}$ with shape $L\times M_{p}\times H\times W$.
-We record the TPRs at all target timestamp as $\boldsymbol{E_{p}}=\{E_{p}\}_{i}^{N_{out}}$.
-Discussion: The time granularity $\delta_{t}$ of TPR exponentially improves with the increase in layers, $L$.
-For an attenuation factor of $r=3$ and a goal to detect motions as brief as $1/1000$ of a second within a $1s$ window ($12\times\Delta t=$ 1), we require only $7$ TPR layers with each layer divided into $2$ intervals ($M_{p}=2$).
+Discussion: The time granularity $\delta_{t}$ of TPR exponentially improves with the increase in layers - $L$.
+For an attenuation factor of $r=3$ and a goal to detect motions as brief as $1/1000$ of a second within a $1s$ window ($2\times\Delta t=1$), we require only $7$ TPR layers with each layer divided into $2$ intervals ($M_{p}=2$).
 Consequently, a TPR with dimensions $7\times 2\times H\times W$ suffices to discern motion down to $1/1000s$.
-Based on the above representation, we obtained $\boldsymbol{E}_{v}$, encapsulating holistic motion, and $\boldsymbol{E}_{p}$, which focuses on regional edges and motion.
+Based on the above representations, we obtained $\bm{E}_{v}$, encapsulating holistic motion, and $\bm{E}_{p}$, which focuses on regional edges and motion.
 3.2 Holistic-Regional Feature Extraction
-This module aims to extract features from regional TPRs $\boldsymbol{E_{p}}$, and the frame $\boldsymbol{I_{in}}$ and the holistic events $\boldsymbol{E_{v}}$ for INR-based spatial-temporal decoding.
-Accordingly, we introduce: (1) the regional event feature extractor, $f_{re}$, for dynamic motion and edges detail capture.
-(2) the holistic event-frame feature extractor, $f_{he}$, for long-term motion dependencies modeling across time and space.
+This module aims to extract features from regional TPRs $\bm{E_{p}}$, and the frame $\bm{I_{in}}$ and the holistic events $\bm{E_{v}}$ for INR-based spatial-temporal decoding.
+Accordingly, we introduce: (1) the regional event feature extractor $f_{re}$, Eq.
+LABEL:eq:feature_extraction (a), for dynamic motion and edges detail capture.
+(2) the holistic event-frame feature extractor $f_{he}$, Eq.
+LABEL:eq:feature_extraction (b), for long-term motion dependencies modeling across time and space.
 Figure 4: Holistic event-frame feature extractor.
@@
 The up-sample module will double the resolution.
-The encoder and decoder have the same structure as Swin-Transformer [37, 35, 16].
-Regional Event Feature Extraction: The input of $f_{re}$ is TPR $E_{p}\in R^{L\times M_{p}\times H\times W}$ in each timestamp $t$.
-Given that $f_{re}$ is invoked $N_{out}$ times, its design needs to strike a balance between efficiency and the ability to model inter-level relationships to ensure that it can capture precise regional motion and edge details effectively.
-Initially, our method applies a convolution layer to increase the dimensions of the feature.
-Then, we use four Swin Transformer Encoder Blocks (STEB) [37, 35, 16] to model the relationship between different pyramid levels.
-The STEB enjoys a large view field and a multi-head attention mechanism that proficiently models distant dependencies and conveys edge information across various pyramid levels, effectively capturing short-term motion.
-Notably, STEB is optimized with fewer parameters, boosting computational efficiency.
-Holistic Event-Frame Feature Extraction: The inputs of $f_{he}$ is $\boldsymbol{I_{in}}$ and $\boldsymbol{E_{v}}$, as shown in Fig.
-3.
+The encoder and decoder have the same structure as Swin-Transformer (Liu et al, 2022b, 2021; Geng et al, 2022).
+[…]
+Figure 5: Temporal embedding.
+Given the input time $t\in[0,1]$, the output is the temporal attention $E_{t}$ derived from a two-layer MLP.
+(b) presents a visualization of the trained $E_{t}$ during $[0,1]$ on real-world dataset (Tulyakov et al, 2022).
+(1) Regional Event Feature Extraction: The regional event feature extractor $f_{re}$ takes as input the Temporal Pyramid Representation (TPR) $\bm{E_{p}}\in\mathbb{R}^{L\times M_{p}\times H\times W}$ at each timestamp $t$, where $L$ denotes the number of TPR layers and $M_{p}$ represents the number of moments within each layer.
+Since $f_{re}$ is invoked $N_{out}$ times, its design must balance computational efficiency with the capacity to model inter-level relationships effectively, ensuring the accurate capture of regional motion and edge details.
+The feature extraction process is divided into two stages: preprocessing and the main feature extraction via Swin Transformer Encoder Blocks (STEB) (Liu et al, 2022b, 2021; Geng et al, 2022).
+In the preprocessing stage, a $1\times 1$ convolutional layer is applied to increase the feature dimensions to $C_{r}$, resulting in a shape of $L\times C_{r}\times H\times W$.
+This step enhances the feature’s capacity to capture intricate details from the event stream.
+In the subsequent feature extraction stage, four STEBs are employed to model the relationships between different pyramid levels.
+STEB utilizes a multi-head self-attention mechanism, which enables the model to effectively capture long-range dependencies and convey edge information across different levels of the pyramid.
+This mechanism improves the model’s ability to capture short-term motion while maintaining computational efficiency through parameter optimization.
+The input and output dimensions of STEB remain consistent throughout the process, preserving the feature shape $L\times C_{r}\times H\times W$.
+Thus, the $f_{re}$ module incorporates both preprocessing and feature extraction steps to ensure efficient extraction of regional motion features while maintaining high computational efficiency.
+Swin Transformer Encoder Blocks (STEB): The input and output shapes of STEB are represented as $L\times C\times H\times W$ for clarity.
+Initially, for a window size of $M\times M$, specifically $4\times 4$ in our implementation, the input is partitioned into disjoint windows of $(M\times M)\times L\times(H/M)\times(W/M)\times C$ dimensions.
+Then, each window is compressed to form a feature map of shape $(M\times M)\times(L\times H\times W/M^{2})\times C$.
+Following this, Layer Normalization (Ba et al, 2016) and window-based multi-head self-attention (Liu et al, 2021; Geng et al, 2022) are computed for each window, succeeded by further transformation via another Layer Normalization and a Multi-Layer Perception.
+Shifted window-based multi-head self-attention (Liu et al, 2021; Geng et al, 2022) is then employed to establish cross-window connections.
+After applying one STEB structure, a second STEB is introduced with an identical configuration, except the input feature window is offset by $(M/2)\times(M/2)$.
+In total, four STEBs are utilized for comprehensive feature extraction in the Regional Event Feature Extraction.
+(2) Holistic Event-Frame Feature Extraction: The inputs of $f_{he}$ is $\bm{I_{in}}$ and $\bm{E_{v}}$, as shown in Fig.
+4.
 First, a convolutional layer processes both frames and events to increase dimensions.
 Motivated by the events feature manifest between successive frames to provide inter-frame motion information.
-We adjusted the first dimension of the event feature to $N_{in}-1$ to account for inter-frame motion losses.
+We construct $N_{in}-1$ event segments so that each segment complements the inter-frame motion between two neighboring RGB frames.
 We then incorporate the STEB to facilitate interactions across varied levels and spatial domains.
@@
 After three iterations, the feature resolution reduces to $1/8$ of its initial size, enlarging the receptive field.
-We then employ a Swin Transformer Decoder Block [37, 35, 16] to fuse features at matching resolutions.
+We then employ a Swin Transformer Decoder Block (Liu et al, 2022b, 2021; Geng et al, 2022) to fuse features at matching resolutions.
 Each of the first three STEBs is followed by an upsampling process, which doubles the resolution while maintaining the channel count.
 Ultimately, this process outputs the feature $F_{g}$.
-For more details, see Suppl.
-Mat..
-Notably, to output $N_{out}$ frames, the holistic event-frame feature extractor $f_{he}$ is called once, capturing the comprehensive feature $F^{g}$.
-Subsequently, for each time $t$, $f_{re}$ extracts regional features $F_{t}^{l}$ from each TPR $E_{p}\in\boldsymbol{E_{p}}$.
-For each regional feature $F_{t}^{l}$, we use addition and Cov1×$\times 1$ 1 operation as fusion function $f_{fu}$ to fuse with holistic feature $F^{g}$ to obtain the output $R_{t}$.
-For each time $t$, the whole process can be described by Eq.
-3, where $E_{p}\in\boldsymbol{E_{p}}$ is the TPR at the specific timestamp $t$.
-$\begin{split}R_{t}=&f_{fu}\left(F^{g},F_{t}^{l}\right);F^{g}=f_{he}\left(\boldsymbol{I_{in}},\boldsymbol{E_{v}}\right);F_{t}^{l}=f_{re}\left(E_{p}\right)\end{split}$ (3)
-Figure 4: Temporal embedding.
-Given the input time $t\in[0,1]$, the output is the temporal attention $E_{t}$ derived from a two-layer MLP.
-(b) presents a visualization of the trained $E_{t}$ during $[0,1]$ on real-world dataset [61].
+[…]
+Figure 6: 7-skip frame interpolation visualization results in Adobe240 dataset (Su et al, 2017).
+Our method (c) more effectively captures the rotating wheels compared to the VideoINR (b), which tends to show noticeable holes and gaps.
+Green circles highlight obvious holes and gaps.
 3.3 INR-based Spatial-Temporal Decoding
@@
 Next, a $1\times 1$ convolution is applied to compress $R_{ts}$ to $C_{ts}$ dimensions, reducing the complexity of the spatial embedding and decoding.
-Spatial Embedding and Decoding: To upscale the temporal embedded features to any desired spatial scale, we utilize a similar approach to previous works [10, 8, 9].
+Spatial Embedding and Decoding: To upscale the temporal embedded features to any desired spatial scale, we utilize a similar approach to previous works (Chen et al, 2022, 2023b, 2021).
 We query the four nearest neighbors in the temporal embedded feature for each spatial coordinate and concatenate these with their distances for spatial embedding.
 A four-layer MLP decoder computes the RGB values, which are then aggregated through area-weighted interpolation for arbitrary-scale super-resolution.
-Similar to [38], we use the Charbonnier loss [30] as the fundamental loss function for training.
-Figure 5: 7-skip frame interpolation visualization results in Adobe240 dataset [56].
-Our method (c) more effectively captures the rotating wheels compared to the VideoINR (b), which tends to show noticeable holes and gaps.
-Green circles highlight obvious holes and gaps.
+Similar to (Lu et al, 2023), we use the Charbonnier loss (Lai et al, 2018) as the fundamental loss function for training.
 4 Experiments
-To facilitate a comprehensive comparison between the frame-base [10, 8] and event-guided methods [38, 62, 61, 23], we employed two simulated datasets [56, 40] and two real-world datasets [61, 51] for our experiments (1) Adobe240 dataset [56] includes 133 videos, each with $720\times 1280$ resolution and a $240fps$ frame rate.
-Following established protocols [10, 8, 70], the dataset is divided into 100 training, 16 validation, and 17 testing subsets.
-We employed the widely-used event simulation method vid2e [15], which accounts for real noise distribution, enhancing our model’s robustness and generalization capabilities.
-(2) GoPro dataset [40] featuring the same resolution and frame-rate with Adobe240 dataset, comprises 22 training and 11 test videos.
-Given its compact size, previous studies [8, 10] have primarily employed the test set for quantitative analysis.
+In this section, we first introduce the datasets, network training and testing details, as well as the evaluation methods in Sec.
+4.1.
+Then, we conduct comparative experiments, which include both joint C-STVSR comparisons and individual comparisons for VFI and VSR tasks in Sec.
+4.2.
+Furthermore, in the following analysis and ablation studies, we explore the performance and characteristics of various components within our framework, as discussed in Sec.
+4.3.
+4.1 Datasets and Implementation Setting
+Table 1: Quantitative metrics (PSNR/SSIM) with 7-skip VFI and $4\times$ VSR. Center Average remain consistent with previous work (Chen et al, 2022, 2023b).
+VFI VSR Params ($M$) GoPro-Center GoPro-Average Adobe-Center Adobe-Average
+Bicubic 19.8 27.04/0.7937 26.06/0.7720 26.09/0.7435 25.29/0.7279
+Su-SloMo
+EDVR 19.8+20.7 28.24/0.8322 26.30/0.7960 27.25/0.7972 25.95/0.7682
+BasicVSR 19.8+6.3 28.23/0.8308 26.36/0.7977 27.28/0.7961 25.94/0.7679
+Bicubic 29.2 26.50/0.7791 25.41/0.7554 25.57/0.7324 24.72/0.7114
+QVI
+EDVR 29.2+20.7 27.43/0.8081 25.55/0.7739 26.40/0.7692 25.09/0.7406
+BasicVSR 29.2+6.3 27.44/0.8070 26.27/0.7955 26.43/0.7682 25.20/0.7421
+Bicubic 24.0 26.92/0.7911 26.11/0.7740 26.01/0.7461 25.40/0.7321
+DAIN
+EDVR 24.0+20.7 28.01/0.8239 26.37/0.7964 27.06/0.7895 26.01/0.7703
+BasicVSR 24.0+6.3 28.00/0.8227 26.46/0.7966 27.07/0.7890 26.23/0.7725
+TimeLens EG-VSR 72.2+2.45 28.85/0.8678 27.54/0.8293 28.11/0.8441 27.42/0.8269
+CBMNet EG-VSR 22.2+2.45 29.22/0.8686 28.51/0.8493 28.28/0.8553 27.89/0.8334
+Zooming Slow Mo 11.10 30.69/0.8847 -/- 30.26/0.8821 -/-
+TMNet 12.26 30.14/0.8692 28.83/0.8514 29.41/0.8524 28.30/0.8354
+Video INR-fixed 11.31 30.73/0.8850 -/- 30.21/0.8805 -/-
+Video INR 11.31 30.26/0.8792 29.41/0.8669 29.92/0.8746 29.27/0.8651
+MoTIF 12.55 31.04/0.8877 30.04/0.8773 30.63/0.8839 29.82/0.8750
+HR-INR (Ours) 8.27 31.97/0.9298 32.13/0.9371 31.26/0.9246 31.11/0.9216
+Table 2: More quantitative comparisons using PSNR/SSIM on the GoPro dataset. Bold indicates the best performance.
+Temporal
+Scale
+Spatial
+Scale
+Su-SloMo
++ LIIF
+DAIN
++ LIIF
+TMNet Video INR MoTIF Ours
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Datasets: To facilitate a comprehensive comparison between the frame-base (Chen et al, 2022, 2023b) and event-guided methods (Lu et al, 2023; Tulyakov et al, 2021, 2022; Jing et al, 2021), we employed two simulated datasets (Su et al, 2017; Nah et al, 2017) and two real-world datasets (Tulyakov et al, 2022; Scheerlinck et al, 2019a) for our experiments.
+(1) Adobe240 Dataset (Su et al, 2017): This dataset comprises 133 videos, each with a resolution of $720\times 1280$ and a frame rate of 240.
+We follow (Chen et al, 2022, 2023b; Xu et al, 2021) to split this dataset into 100 training, 16 validation, and 17 test sets.
+We employed the widely-used event simulation method vid2e (Gehrig et al, 2020), which accounts for real noise distribution, ensuring robust neural network training with enhanced generalization.
+In generating the inputs and ground truth of the network, we adopted and extended the previous works (Chen et al, 2023b, 2022) to accommodate multi-frame input.
+Specifically, input and ground truth (GT) frames are selected via sliding windows.
+We define the window size as $W$, the number of input frames as $N_{in}$, and the interval as $S$.
+They interrelate as: $W=(N_{in}-1)*(S+1)+1$ For instance, with 4 frames input at 7-frame intervals, 25 frames are chosen.
+The 1st, 9th, 17th, and 25th frames become the input after down-sampling, while frames 1-25 serve as the GT.
+We adopted two strategies in line with prior works (Chen et al, 2022, 2023b): I) A fixed magnification set at $4\times$ the input resolution, and II) A variable enlargement strategy, wherein the scaling factor is governed by a $\mathcal{U}(1,8)$ distribution.
+(2) GoPro dataset (Nah et al, 2017) featuring the same resolution and frame-rate with  Adobe240 dataset, comprises 22 training and 11 test videos.
+Given its compact size, previous studies (Chen et al, 2023b, 2022) have primarily employed the test set for quantitative analysis.
 We follow this to be consistent with established practice.
-For datasets (1) and (2) during 7-skip frame interpolation, we input 4 frames where each pair of adjacent frames is separated by 7 frames, allowing a total of 25 frames to serve as GTs.
-(3) BS-ERGB [61] recorded using a spectroscope, includes real-world paired events and frames.
-Previous studies, e.g., TimeLens++[61], initially pre-trained on the Vime90k simulation dataset[72] before fine-tuning on BS-ERGB.
-In contrast, we opted to pre-train our model on the Adobe240 dataset before fine-tuning.
-Notably, the Vime90k dataset is larger in scale than the Adobe240 dataset.
-During fine-tuning, we also used perceptual loss [25] with weight $0.1$ to be consistent with previous methods [62, 61] for fair comparison.
-(4) CED [51] is a real-world dataset in VSR.
-To fairly compare the previous research [23, 38], only this data set is used for training for VSR comparison.
-Implementation Details: Our model is trained using Pytorch [48], employing the Adam optimizer [29].
-Referring to the VideoINR [10], our training consists of two stages.
-(1) Train frame interpolation under a fixed spatial up-sampling (4×4$\times$4 ×), over 70 epochs, starting with a learning rate of $45e-$ 4.
-(2) Train frame interpolation under random space upsampling rate in $\mathcal{U}(1,8)$, spanning 30 epochs with the learning rate $55e-$ 5.
+(3) BS-ERGB (Tulyakov et al, 2022) recorded using a spectroscope, includes real-world paired events and frames with a $970\times 625$ resolution at 28 fps.
+Of the 123 videos in the dataset, 47 are allocated for training, 19 for validation, and 26 for testing.
+Each video contains between 100 to 600 frames.
+We opted to pre-train our model on the Adobe240 dataset before fine-tuning.
+During fine-tuning, we also used perceptual loss (Johnson et al, 2016) with weight $0.1$ to be consistent with previous methods (Tulyakov et al, 2021, 2022) for fair comparison.
+(4) CED (Scheerlinck et al, 2019a) is a real-world dataset in VSR.
+To fairly compare the previous research (Jing et al, 2021; Lu et al, 2023), only this data set is used for training for VSR comparison.
+It is worth noting that the dataset ALPIX-VSR (Lu et al, 2023) can also be used for event-guided VSR; however, its events are not stream-based  (Alpsentek, 2024), making it unsuitable for events TPR.
+Therefore, we use CED as the primary dataset for comparison.
+Figure 7: Timestamps and corresponding PSNR for each frame during $4\times$ VSR and 7-skip VFI on the Adobe240 dataset (Su et al, 2017).
+Implementation Details: Our model is trained using Pytorch (Paszke et al, 2019), employing the Adam optimizer (Kingma and Ba, 2014).
+Referring to the VideoINR (Chen et al, 2022), our training consists of two stages.
+(1) Train frame interpolation under a fixed spatial up-sampling ($4\times$), over 70 epochs, starting with a learning rate of 5e45e-4.
+(2) Train frame interpolation under random space upsampling rate in $\mathcal{U}(1,8)$, spanning 30 epochs with the learning rate 5e55e-5.
 We randomly choose 20 frames from a pool of 25 frames as ground truth.
 Data augmentation is implemented via Random Crop, extracting $512\times 512$ areas from frames and events, with the input resolution dynamically determined by a randomly chosen upsampling ratio $s$.
-To optimize memory usage and accelerate speed, we implemented the mixed precision strategy [39, 12].
-All experiments are performed on an NVIDIA A800 computing card.
-Please refer to the Suppl.
-Mat.
-for more details.
-Evaluation: To ensure the fairness, we adopted PSNR [78], SSIM [67], and LPIPS [78] as quantitative evaluation metrics.
-Aligning with prior works [10, 8] for consistency, we use only the $Y$-channel for GoPro and Adobe datasets and all three RGB channels for BS-ERGB and CED datasets.
-Table 1: Quantitative metrics (PSNR/SSIM) with 7-skip VFI and 4×4$\times$4 × VSR. Center Average remain consistent with previous work [10, 8].
-VFI VSR Params ($M$) GoPro-Center GoPro-Average Adobe-Center Adobe-Average
-Bicubic 19.8 27.04/0.7937 26.06/0.7720 26.09/0.7435 25.29/0.7279
-Su-SloMo [22] EDVR [66] 19.8+20.7 28.24/0.8322 26.30/0.7960 27.25/0.7972 25.95/0.7682
-BasicVSR [5] 19.8+6.3 28.23/0.8308 26.36/0.7977 27.28/0.7961 25.94/0.7679
-[…]
-QVI [71] EDVR [66] 29.2+20.7 27.43/0.8081 25.55/0.7739 26.40/0.7692 25.09/0.7406
-BasicVSR [5] 29.2+6.3 27.44/0.8070 26.27/0.7955 26.43/0.7682 25.20/0.7421
-[…]
-DAIN [3] EDVR [66] 24.0+20.7 28.01/0.8239 26.37/0.7964 27.06/0.7895 26.01/0.7703
-BasicVSR [5] 24.0+6.3 28.00/0.8227 26.46/0.7966 27.07/0.7890 26.23/0.7725
-Zooming Slow Mo [69] 11.10 30.69/0.8847 -/- 30.26/0.8821 -/-
-TMNet [70] 12.26 30.14/0.8692 28.83/0.8514 29.41/0.8524 28.30/0.8354
-Video INR-fixed [10] 11.31 30.73/0.8850 -/- 30.21/0.8805 -/-
-Video INR [10] 11.31 30.26/0.8792 29.41/0.8669 29.92/0.8746 29.27/0.8651
-MoTIF [8] 12.55 31.04/0.8877 30.04/0.8773 30.63/0.8839 29.82/0.8750
-HR-INR (Ours) 8.27 31.97/0.9298 32.13/0.9371 31.26/0.9246 31.11/0.9216
-Table 2: More quantitative comparisons using PSNR/SSIM on the GoPro dataset.
-Temporal Scale Spatial Scale Su-SloMo [22] + LIIF [9] DAIN [3] + LIIF [9] TMNet [70] Video INR [10] MoTIF [8] Ours
-12$\times$ 4$\times$ 25.07/0.7491 25.14/0.7497 26.38/0.7931 27.32/0.8141 27.77/0.8230 28.87/0.8854
-12$\times$ 6$\times$ 22.91/0.6783 22.92/0.6785 - 24.68/0.7358 26.78/0.7908 27.14/0.8173
-16$\times$ 4$\times$ 24.42/0.7296 24.20/0.7244 24.72/0.7526 25.81/0.7739 25.98/0.7758 27.29/0.8556
-16$\times$ 6$\times$ 23.28/0.6883 $2$ 23.86/0.7123 25.34/0.7527 26.09/0.7954
-6$\times$ 1$\times$ - - - 32.34/0.9545 34.77/0.9696 38.53/0.9735
+To optimize memory usage and accelerate speed, we implemented the mixed precision strategy (Micikevicius et al, 2017; Das et al, 2018).
+All experiments are performed on an NVIDIA A800.
+Figure 8: 3-skip frame interpolation visualization results in BS-ERGB dataset (Tulyakov et al, 2022).
+Our method accurately captures local non-linear motion (e.g., balls in (1.d) and (2.d)), surpassing TimeLens (Tulyakov et al, 2022), which exhibits ghosting and holes (green circles).
+A yellow arrow shows TimeLens’s inaccurate ball positioning.
+Evaluation: To ensure the fairness, we adopted PSNR (Zhang et al, 2018), SSIM (Wang et al, 2004), and LPIPS (Zhang et al, 2018) as quantitative evaluation metrics.
+For a fair comparison on GoPro and Adobe datasets, we follow previous work (Chen et al, 2022, 2023b) by using only the $Y$-channel, while for the BS-ERGB and CED datasets, we employ all three RGB channels.
 4.2 Comparison Experiments
-Space-time Super-resolution: We conduct space-time super-resolution comparison experiments on the Adobe240 and GoPro datasets.
+Since our method is the first event-based C-STVSR approach, we compare it with both frame-based methods and cascade methods for C-STVSR, as discussed in Sec.
+4.2.1.
+Additionally, our method also demonstrates superior performance in individual tasks, VFI and VSR.
+The comparisons for these individual tasks are discussed in Sec.
+4.2.2.
+Figure 9: Visualization results on a real-world dataset in a large object long-distance motion scene.
+(a) and (b) represent the input frames, (c) shows the corresponding events, (d) is the result of the TimeLens method (TL), (e) shows the interpolation by CBMNet (CL), (f) presents the result of our method, and (j) is the ground truth (GT), and (h) displays the local events at each time step.
+4.2.1 Space-time Super-resolution
+We conduct space-time super-resolution comparison experiments on the Adobe240 and GoPro datasets.
 We categorized the comparison methods into three groups.
-(I) Frame-based cascade methods: VFI methods, e.g., Super SloMo [22] and DAIN [3] followed by VSR methods, e.g., EDVR [66] and BasicVSR [5].
-(II) Fixed STVSR methods: e.g., Zooming Slow-Mo [69] and TMNet [70].
-(III) Frame-based C-STVSR methods: VideoINR [10] and MoTIF [8].
+(I) Cascade methods: VFI methods, e.g., Super SloMo (Jiang et al, 2018), DAIN (Bao et al, 2019a) and TimeLens (Tulyakov et al, 2021) followed by VSR methods, e.g., EDVR (Wang et al, 2019), BasicVSR (Chan et al, 2021) and EG-VSR (Lu et al, 2023).
+(II) Fixed STVSR methods: e.g., Zooming Slow-Mo (Xiang et al, 2020) and TMNet (Xu et al, 2021).
+(III) Frame-based C-STVSR methods: VideoINR (Chen et al, 2022) and MoTIF (Chen et al, 2023b).
 The numerical comparison is presented in Tab.
 1.
-It is evident that C-STVSR methods consistently outperform cascade and fixed STVSR methods.Our method achieves the highest performance in both datasets with the smallest model size.
-In the GoPro dataset, our method improves the center frame by 0.93 $dB$ and 0.0415 SSIM, and on average by 2.09 $dB$ and 0.0598 SSIM compared to the best method, MoTIF [8].
-Similarly, in the Adobe240 dataset, our method outperforms MoTIF [8] by 0.63 $dB$ and 0.0407 SSIM for the -center, and on -average by 1.29 $dB$ and 0.0466 SSIM.
+It is evident that C-STVSR methods consistently outperform cascade and fixed STVSR methods.
+Our method achieves the highest performance in both datasets with the smallest model size.
+In the GoPro dataset, our method improves the center frame by 0.93 $dB$ and 0.0415 SSIM, and on average by 2.09 $dB$ and 0.0598 SSIM compared to the best method, MoTIF (Chen et al, 2023b).
+Similarly, in the Adobe240 dataset, our method outperforms MoTIF (Chen et al, 2023b) by 0.63 $dB$ and 0.0407 SSIM for the -center, and on -average by 1.29 $dB$ and 0.0466 SSIM.
 It is worth noting that the difference between our method and other methods is more pronounced for the -average frames than the -center frame.
-This observation suggests that our method demonstrates enhanced temporal stability and superior adaptability across varying interpolation intervals, an advantage not shared by previous methods [10, 8].
-For more analysis, please refer to Sec.
-4.2 and Fig.
-8 in appendix.
-Table 3: Temporal super-resolution results, i.e.,, VFI, on BS-ERGB dataset [61].
+This observation suggests that our method demonstrates enhanced temporal stability, as shown in Fig.
+7, and superior adaptability across varying interpolation intervals, an advantage not shared by previous methods (Chen et al, 2022, 2023b).
+Tab.
+2 presents additional comparative experiments for arbitrary spatial and temporal super-resolution.
+Our method consistently outperforms other methods, even in extreme space-time upsample scales, such as $16\times$ temporal upscale and $6\times$ spatial upscale.
+The visualization results, Fig.
+1 (c) and Fig.
+6, also demonstrate that our method effectively models regional nonlinear motion, e.g., the wheel rotation — a capability not achieved by previous method VideoINR (Chen et al, 2022).
+For more video visualization results, please refer to the Supplementary Materials.
+Stability Across Temporal Sequences: Our method demonstrates temporal stability, accurately estimating motion states at each time point during frame interpolation, as shown in Fig.
+7.
+This is the key reason why our method outperforms previous approaches.
+Specifically, whether in the GoPro-Average and Adobe-Average of Tab.
+1, or the 12-skip and 16-skip tests of Tab.
+2, our method significantly outperforms previous methods by at least $1.2dB$ with $4\times$ VSR.
+Figure 10: Visualization results on a real-world dataset in a small object long-distance motion scene.
+(a) and (b) represent the input frames, (c) shows the corresponding events, (d) is the result of the TimeLens method (TL), (e) shows the interpolation by CBMNet-Large (CL), (f) presents the result of our method, and (j) is the ground truth (GT), and (h) displays the local events at each time step.
+Table 3: Temporal super-resolution results, i.e.,, VFI, on BS-ERGB dataset (Tulyakov et al, 2022).
 1-skip 3-skip
 Methods Params ($M$) Event PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
-FLAVR [26] - 25.95 - 0.086 20.90 - 0.151
-DAIN [3] 24.0 25.20 - 0.067 21.40 - 0.113
-Super SloMo [22] 19.8 - - - 22.48 - 0.115
-QVI [71] 29.2 - - - 23.20 - 0.110
-TimeLens [62] 72.2 28.36 - 0.026 27.58 - 0.031
-TimeLens++ [61] 53.9 28.56 - 0.022 27.63 - 0.026
-CBMNet [28] 15.4 29.32 0.815 - 28.46 0.806 -
-CBMNet-Large [28] 22.2 29.43 0.816 - 28.59 0.808 -
+FLAVR (Kalluri et al, 2023) - 25.95 - 0.086 20.90 - 0.151
+DAIN (Bao et al, 2019a) 24.0 25.20 - 0.067 21.40 - 0.113
+Super SloMo (Jiang et al, 2018) 19.8 - - - 22.48 - 0.115
+QVI (Xu et al, 2019) 29.2 - - - 23.20 - 0.110
+TimeLens (Tulyakov et al, 2021) 72.2 28.36 - 0.026 27.58 - 0.031
+TimeLens++ (Tulyakov et al, 2022) 53.9 28.56 - 0.022 27.63 - 0.026
+CBMNet (Kim et al, 2023) 15.4 29.32 0.815 - 28.46 0.806 -
+CBMNet-Large (Kim et al, 2023) 22.2 29.43 0.816 - 28.59 0.808 -
 HR-INR (Our) 8.3 29.66 0.828 0.011 28.59 0.814 0.021
-Tab.
-2 presents additional comparative experiments for arbitrary spatial and temporal super-resolution.
-Our method consistently outperforms other methods, even in extreme space-time upsample scales, such as 16×16$\times$16 × temporal upscale and 6×6$\times$6 × spatial upscale.
-The visualization results, Fig.
-1 (c) and Fig.
-5, also demonstrate that our method effectively models regional nonlinear motion, e.g., the wheel rotation — a capability not achieved by previous method VideoINR [10].
-For more visualization results, please refer to the Suppl.
-Mat.
-Figure 6: 3-skip frame interpolation visualization results in BS-ERGB dataset [61].
-Our method accurately captures local non-linear motion (e.g., balls in (1.d) and (2.d)), surpassing TimeLens [61], which exhibits ghosting and holes (green circles).
-A yellow arrow shows TimeLens’s inaccurate ball positioning.
-Separate Comparison of Event-based VFI and VSR: We also compared our method with previous approaches in separate VFI [61, 62, 28] and VSR [38, 23] tasks.
+Fig.
+7 shows the relationship between timestamps and PSNR, validating the greater stability of our method compared to VideoINR  (Chen et al, 2022), especially around the $0.5s$ mark, where VideoINR experiences a notable decrease.
+This observation is also reflected in the visualization results of Fig.
+16 and the following Fig.
+8910.
+Whether using real-world or simulated data, our method demonstrates superior temporal consistency, exhibiting fewer artifacts compared to previous methods such as VideoINR (Chen et al, 2022), TimeLens (Tulyakov et al, 2021) and CMB-Net (Kim et al, 2023).
+This also highlights the real-world effectiveness of our method’s temporal stability.
+The main reason for this is that previous methods (Chen et al, 2022, 2023b; Tulyakov et al, 2021) suffer from instability in motion estimation at timestamps far from the reference frames (at 0 and 11 timestamp), which negatively impacts shape and edge estimation for frame interpolation.
+In contrast, our method benefits from the intermediate motion details provided by events TPR and the model’s ability to capture long temporal sequences, enabling it to maintain higher stability across the temporal sequence and deliver superior performance.
+Figure 11: $4\times$ video super-resolution visualization results in CED dataset (Scheerlinck et al, 2019b).
+Table 4: Spatial super-resolution results on CED (Scheerlinck et al, 2019b). * denotes values from pre-trained models.
+$4\times$ $2\times$
+Methods Params ($M$) Events PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
+TDAN (Tian et al, 2020) 1.97 27.88 0.8231 33.74 0.9398
+SOF (Wang et al, 2020) 1.00 27.00 0.8050 31.84 0.9226
+RBPN (Haris et al, 2019) 12.18 29.80 0.8975 36.66 0.9754
+VideoINR* (Chen et al, 2022) 11.31 25.53 0.7871 26.77 0.7938
+E-VSR (Jing et al, 2021) 412.42 30.15 0.9052 37.32 0.9783
+EG-VSR (Lu et al, 2023) 2.45 31.12 0.9211 38.69 0.9771
+HR-INR (Our) 8.27 32.15 0.9658 42.01 0.9905
+4.2.2 Separate Comparison of Event-based VFI and VSR
+We also compared our method with previous approaches in separate VFI (Tulyakov et al, 2022, 2021; Kim et al, 2023) and VSR (Lu et al, 2023; Jing et al, 2021) tasks.
 The results can be seen in Tab.
@@
 4 respectively.
-In the VFI task, our method surpasses TimeLens++ [61] by 1.1 $dB$ for 1-skip and 0.17 $dB$ for 3-skip VFI, and CBMNet [28] by 0.23 $dB$ for 1-skip.
-Additionally, our model’s size is merely $1/7$ and $1/3$ that of TimeLens++ [61] and CBMNet [28], respectively.
+Comparison with Event-based VFI: In the VFI task, our method surpasses TimeLens++ (Tulyakov et al, 2022) by 1.1 $dB$ for 1-skip and 0.17 $dB$ for 3-skip VFI, and CBMNet (Kim et al, 2023) by 0.23 $dB$ for 1-skip in Tab.
+3.
+Additionally, our model’s size is merely $1/7$ and $1/3$ that of TimeLens++ (Tulyakov et al, 2022) and CBMNet (Kim et al, 2023), respectively.
 In the visualization aspect, our approach excels in modeling non-linear motion and long-term dependencies, as evident in Fig.
-6 in appendix.
-It precisely forecasts the positions of small balls (Fig.
+8, where the subject is juggling, the motion of small, fast-moving objects (balls) is challenging to capture.
+Our method precisely forecasts the positions of small balls (Fig.
 8 (1.d)) and the soccer ball (Fig.
-6 (2.d)) at intermediate timestamps, outperforming previous methods.
-For the VSR task, our method outperforms EG-VSR [38] by 1.03 $dB$ and 0.447 SSIM for 4 $\times$ super-resolution, and by 3.02 $dB$ and 0.0134 SSIM for 2 $\times$ super-resolution.
+8 (2.d)) at intermediate timestamps.
+Competing methods exhibit noticeable ghosting and fail to preserve the spatial consistency of the balls.
+This demonstrates our model’s ability to handle high-speed motion effectively by leveraging its dual-branch feature extraction.
+Table 5: Ablation studies in Adobe-Average (Su et al, 2017) ($4\times$ and 7-skip). The †  symbol marks the line for comparison with other lines.
+Case Events TPR
+Temporal
+Embedding
+Temporal
+Dim ($C_{t}$)
+Input
+Frames
+PSNR $\uparrow$
+SSIM $\uparrow$
+Case#1 Learning 640 4 26.84 ( 4.27-4.27) 0.8366 ( 0.0850-0.0850)
+Case#2 Learning 640 4 29.69 ( 1.41-1.41) 0.8974 ( 0.0242-0.0242)
+Case#3 † Learning 640 4 31.11 0.9216
+Case#4 Sinusoid 640 4 30.42 ( 0.69-0.69) 0.9120 ( 0.0096-0.0096)
+Case#5 Learning 320 4 28.44 ( 2.67-2.67) 0.8700 ( 0.0516-0.0516)
+Case#6 Learning 640 2 30.41 ( 0.70-0.70) 0.9151 ( 0.0065-0.0065)
+Case#7 Learning 640 3 30.72 ( 0.39-0.39) 0.9174 ( 0.0042-0.0042)-
+Furthermore, we also explore the performance of our method in scenarios involving motions of objects of different sizes.
 Fig.
-7 demonstrates our model’s effective VSR performance on the real-dataset CED, highlighting its ability to reproduce sharp edges and robustness to noise.
-For more visualization results, see the Suppl.
-Mat.
-Table 4: Spatial super-resolution results on CED dataset. * denotes the values obtained from the official pre-trained models.
-      4×4$\times$4 ×       2×2$\times$2 ×
-      Methods       Params ($M$)       Events       PSNR $\uparrow$       SSIM $\uparrow$       PSNR $\uparrow$       SSIM $\uparrow$
-      DUF [24]       1.90       ✗       24.43       0.8177       31.83       0.9183
-      TDAN [60]       1.97       ✗       27.88       0.8231       33.74       0.9398
-      SOF [64]       1.00       ✗       27.00       0.8050       31.84       0.9226
-      RBPN [19]       12.18       ✗       29.80       0.8975       36.66       0.9754
-      BasicVSR++ [6]       7.30       ✗       14.76       0.1641       -       -
-      VideoINR* [10]       11.31       ✗       25.53       0.7871       26.77       0.7938
-      E-VSR [23]       412.42       ✓       30.15       0.9052       37.32       0.9783
-      EG-VSR [38]       2.45       ✓       31.12       0.9211       38.69       0.9771
-      HR-INR (Our)       8.27       ✓       32.15       0.9658       42.01       0.9905
-Figure 7: 4×4$\times$4 × video super-resolution visualization results in CED dataset [52].
-Table 5: Ablation studies in Adobe-Average [56] (4×4$\times$4 × and 7-skip). The †  symbol marks the line for comparison with other lines.
-Case Events TPR Temporal Embedding Temporal Dim ($C_{t}$) Input Frames PSNR $\uparrow$ SSIM $\uparrow$
-Case#1 Learning 640 4 26.84 ( $-4.27$) 0.8366 ( $-0.0850$)
-Case#2 Learning 640 4 29.69 ( $-1.41$) 0.8974 ( $-0.0242$)
-Case#3 † Learning 640 4 31.11 0.9216
-Case#4 Sinusoid 640 4 30.42 ( $-0.69$) 0.9120 ( $-0.0096$)
-Case#5 Learning 320 4 28.44 ( $-2.67$) 0.8700 ( $-0.0516$)
-Case#6 Learning 640 2 30.41 ( $-0.70$) 0.9151 ( $-0.0065$)
-Case#7 Learning 640 3 30.72 ( $-0.39$) 0.9174 ( $-0.0042$)-
-Table 6: Ablation studies for TPR levels and moments in Adobe-Average [56] (4×4$\times$4 × and 7-skip). The †  symbol marks the line for comparison with other lines.
-      Case       TPR Level ($L$)       TPR Moments ($M_{p}$)       PSNR $\uparrow$       SSIM $\uparrow$
-      Case#1       3       3       29.93 ( $-1.18$)       0.9011 ( $-0.0205$)
-      Case#2       5       3       30.32 ( $-0.79$)       0.9165 ( $-0.0051$)
-      Case#3       7       3       30.78 ( $-0.33$)       0.9187 ( $-0.0029$)
-      Case#4 †       7       9       31.11       0.9216
-      Case#5       7       18       31.18 ( $+0.07$)       0.9228 ( $+0.0012$)
-Table 7: Ablation in BS-ERGB (TimeLens++) dataset [61].
+9, 10 demonstrate the superiority of our method in recovering large-scale and long-distance motions across various real-world scenarios.
+These results highlight the effectiveness of our approach in fusing both regional and holistic information to handle complex motion dynamics, while simultaneously reducing ghosting artifacts that are commonly seen in competing methods.
+Fig.
+9 shows large object long-distance motion, such as the horse and rider jumping, our method shows clear advantages.
+The silhouette of the horse and the rider’s position are reconstructed with remarkable clarity and consistency.
+In contrast, competing methods introduce significant blurring and fail to preserve the integrity of the subject’s shape, revealing their limitations in handling large-scale motion over time.
+Fig.
+10 shows small object long-distance motion.
+Our method excels in reconstructing fine-grained motion details, such as the movement of the horse’s legs.
+Unlike CMB-Net and TimeLens, which either fail to capture these intricate details or introduce substantial motion blur, our approach accurately restores the distinct positions of the horse’s hooves.
+This ability arises from the integration of local motion features captured by the regional branch, along with the long-term temporal dependencies modeled by the holistic branch.
+Further analysis can be found in Sec.
+4.3.
+Across all scenarios, our method generates outputs with fewer ghosting artifacts.
+The combination of regional and holistic features allows our approach to capture both short-term and long-term motion dependencies, resulting in more realistic and temporally coherent video outputs.
+These results validate the robustness and generalizability of our method across diverse motion patterns and scales.
+Comparison with Event-based VSR: For the VSR task, our method outperforms EG-VSR (Lu et al, 2023) by 1.03 $dB$ and 0.447 SSIM for 4 $\times$ super-resolution, and by 3.02 $dB$ and 0.0134 SSIM for 2 $\times$ super-resolution.
+As shown in Fig.
+10, our method produces visually sharper results compared to EG-VSR and other baseline methods.
+In Fig.
+11, we present video super-resolution visualization results on the CED dataset (Scheerlinck et al, 2019b), where our approach effectively restores fine details, such as the text on the building and edges around structures, which are blurred or distorted in competing methods.
+The corresponding error map (Fig.
+11 (g)) further demonstrates our method’s reduced errors, especially around high-contrast regions.
+Figure 12: Feature visualization on real data (Tulyakov et al, 2022): (f) shows the holistic feature, $F^{g}$, derived from multiple frames and events; (g) depict the regional features ($F_{t}^{g}$), highlighting the capability to capture local motion.
+4.3 Ablation and Analytical Studies
+Ablation and analytical studies conducted on the Adobe240 (Su et al, 2017) and BS-ERGB (Tulyakov et al, 2022) datasets unveiled several critical insights.
+On the Adobe240 dataset, we executed simultaneous 7-skip VFI and $4\times$ VSR tests, as shown in Tab.
+5 and Tab.
+6, while on the BS-ERGB dataset, 1-skip and 3-skip VFI were performed in Tab.
+7.
+Figure 13: Visualization of Temporal Embedding.
+Figure (a) shows the original visualization of Temporal Embedding, while figure (b) displays the results after sorting.
+The sorting is based on the size of the first element, arranged in ascending order.
+We use MLP decoding, where the order is not crucial.
+However, to more clearly demonstrate the outcomes of Temporal Embedding learning, we have chosen to present the sorted results.
+Table 6: Ablation studies for TPR levels and moments in Adobe-Average Su et al (2017) ($4\times$ and 7-skip).
+The †  symbol marks the line for comparison with other lines.
+”Captured Moment” refers to the temporal resolution of the last layer of the TPR, which is calculated by Eq.
+4.
+Case TPR Level ($L$) TPR Moments ($M_{p}$) Captured Moment PSNR $\uparrow$ SSIM $\uparrow$
+Case#1 3 3 1 / 81 29.93 ( 1.18-1.18) 0.9011 ( 0.0205-0.0205)
+Case#2 5 3 1 / 729 30.32 ( 0.79-0.79) 0.9165 ( 0.0051-0.0051)
+Case#3 7 3 1 / 6561 30.78 ( 0.33-0.33) 0.9187 ( 0.0029-0.0029)
+Case#4 † 7 9 1 / 19683 31.11 0.9216
+Case#5 7 18 1 / 39366 31.18 ( +0.07+0.07) 0.9228 ( +0.0012+0.0012)
+Table 7: Ablation in BS-ERGB (TimeLens++) dataset Tulyakov et al (2022).
 1-skip 3-skip
@@
 28.25 0.8187 0.018 26.65 0.7867 0.039
-29.66 ( $+1.41$) 0.8281 ( $+0.0094$) 0.011 ( $-0.007$) 28.59 ( $+1.94$) 0.8140 ( $+0.0273$) 0.021 ( $-0.018$)
-4.2 Ablation and Analytical Studies
-Ablation and analytical studies conducted on the Adobe240 [56] and BS-ERGB [61] datasets unveiled several critical insights.
-On the Adobe240 dataset, we executed simultaneous 7-skip VFI and 4×4$\times$4 × VSR tests, as shown in Tab.
-5 and Tab.
-6, while on the BS-ERGB dataset, 1-skip and 3-skip VFI were performed in Tab.
-7.
+29.66 ( +1.41+1.41) 0.8281 ( +0.0094+0.0094) 0.011 ( 0.007-0.007) 28.59 ( +1.94+1.94) 0.8140 ( +0.0273+0.0273) 0.021 ( 0.018-0.018)
 Events Gain: Tab.
@@
 Adding events alone substantially raised PSNR by $2.85dB$ and SSIM by $0.06$.
-Event TPR Enhancement: Incorporating event TPR enhances performance, with PSNR and SSIM increasing with layer count, peaking at seven layers, as shown in Tab.
-5-Case#2 and Tab.
+Event TPR Analysis: Incorporating event TPR significantly enhances performance, with both PSNR and SSIM improving as the layer count increases, peaking at seven layers, as shown in Tab.
+5 and Tab.
 6.
-However, expanding ETPR moments from three to seven only improves PSNR by $0.33dB$, with minimal gains beyond that, suggesting a plateau in performance benefits.
-The TPR also demonstrates enhancement on the BS-ERGB dataset, as shown in Tab.
-7, yielding an increase of $0.92dB$ for 1-skip and an improvement of $1.13dB$ for 3-skip.
-This indicates that the benefits of TPR become more pronounced with the increase in the number of skips.
-Time Embedding Method: Table 5-Case#3#4 shows Sinusoid embedding’s results.
-It’s outperformed by learning-based methods, confirmed by previous research [59, 50, 82, 1], due to their superior high-frequency information capture.
+Specifically, in Tab.
+6, as the TRP Level $L$ increases, the moments captured by the TPR become more precise.
+For example, when the TRP Level rises from 3 in Tab.
+6-Case#1, to 7 in Tab.
+6-Case#4, the PSNR shows an increase of approximately 1.18.
+Moreover, the model’s performance further improves with the increase in TPR Moments $M_{p}$.
+However, it is important to note that when both $L$ and $M_{p}$ are relatively high, the performance improvement tends to plateau.
+Additionally, the TPR enhancement is also evident on the BS-ERGB dataset, as shown in Tab.
+7, where it yields an increase of $0.92$ dB for 1-skip and an improvement of $1.13$ dB for 3-skip.
+Analysis of Event TPR in Visualization: Building upon the performance analysis of TPR, we now examine its effectiveness in visualization.
+As shown in Fig.
+12, TPR’s regional features emphasize localized motion, demonstrating its ability to capture fine-grained, short-term dynamics.
+This capability highlights TPR as a novel event representation for millisecond-scale motion at high temporal precision.
+For instance, the regional features (Fig.12(h)) focus on localized, short-term motion, while the holistic features (Fig.12(d)) capture the broader, long-term context.
+Together, these two types of features allow our method to effectively separate dynamic motion from static background information.
+In the Global Motion Scene, as shown in Fig.
+12, where vehicles move against a stationary urban background, the regional features capture the intricate, millisecond-level motion of the vehicles, while the holistic features retain the overall structure of the scene, including static elements such as traffic signs and buildings.
+This dual representation enables our method to handle complex dynamic scenes seamlessly, where fast-moving objects coexist with stationary elements.
+In the Local Motion Scene, as shown in Fig.
+12, the visualization of a basketball player dribbling the ball reveals that the holistic features primarily encode static background details like trees and stationary objects, while the regional features focus on the fast, dynamic motion of the basketball.
+This demonstrates TPR’s strength in isolating short-term motion, even for high-speed moving objects.
+These visualizations provide strong evidence of TPR’s capability to capture fine-grained, short-term motion through its hierarchical structure.
+By focusing on temporal granularity, event data complements frame-based information and enables the extraction of rich local motion features.
+We also encourage readers to refer to the supplementary videos for further demonstrations, where these cases are presented in more detail, showcasing how TPR integrates both event and frame data to effectively handle complex motion patterns.
+This visualization underscores the unique advantages of TPR, combining regional and holistic information to enable robust spatiotemporal super-resolution.
+Time Embedding Method: Tab.
+5-Case#3#4 presents the results for sinusoid embeddings.
+The learning-based methods perform better due to their superior ability to capture high-frequency information (Ramasinghe and Lucey, 2023; Attal et al, 2022).
+This advantage is further demonstrated in Fig.
+13, which visualizes the time embedding features.
+The learning-based embeddings not only effectively capture periodic positional representations but also offer a more comprehensive expression of exposure information.
 Temporal Dimension Impact: The INR temporal dimension significantly influences performance.
@@
 This highlights the need to balance dimensionality and training stability.
-Input Frames: Tab.
+Figure 14: Visualization of Holistic Features for different frames as input.
+Analysis of Input Frames and Long-distance Modeling: Tab.
 5-Case#3#6#7 illustrates the impact of varying input frame counts on the final results.
@@
 This indicates a clear advantage of multi-frame inputs in modeling longer-term dependencies.
-Moreover, even with two frames, our method also outperforms previous works [10, 8].
-5 Conclusion
-Our work introduced the first event-guided continuous space-time video super-resolution method.
-The main contributions are: (I) Event temporal pyramid representation for capturing short-term dynamic motion; (II) A feature extraction process combining holistic and regional features to manage motion dependencies; (III) A spatiotemporal decoding process based on implicit neural representation, avoiding traditional optical flow and achieving stable frame interpolation through temporal-spatial embedding.
-Limitations and Future Work: Training a robust C-STVSR network requires extensive time and computational resources.
-Addressing these challenges remains an important area for future research.
-Due to space limitations, please refer more details, analysis and visualization results in the appendices.
-Appendix Overview
-The appendix of this paper consists of five sections:
-  1. 1.
-    (Sec. A) More Details about the Event Generation Model:
-    This section introduces the event generation model, providing insights into why events are effective for C-STVSR.
-  2. 2.
-    (Sec. B) More Details about the Network Structure:
-    Detailed information on the network architecture is provided, including comprehensive descriptions of the inputs, outputs, and intermediate processes of each module.
-  3. 3.
-    (Sec. C) More Details about the Experimental Settings:
-    This section describes additional experimental settings to ensure the accurate replication of our method.
-  4. 4.
-    (Sec. D) Additional Experimental Results and Analysis :
-    Further experimental analysis and results are offered, demonstrating the superiority of our approach.
-  5. 5.
-    (Sec. E) Additional Visualization Results:
-    This section presents more visual materials, including images and videos.
-Appendix A More Details about Event Generation Model
-Events are discrete points capturing the positive and negative changes in pixel brightness.
-Their generation hinges on brightness alterations within the logarithmic domain.
-Specifically, an event point $e=(x,y,t,p)$ is triggered and logged upon meeting certain criteria.
-Suppose $L(x,y,t)$ represents the brightness at point $(x,y)$ at any given time $t$.
-The event is recorded if the absolute difference $\Delta L=log\left(L(x,y,t)\right)-L\left(x,y,t-\Delta t\right)$ surpasses a predetermined threshold $C$, as formulated as Eq.
-4.
-$p=\left\{\begin{aligned} +1,&\Delta L>C\\ -1,&\Delta L<-C\end{aligned}\right.$ (4)
-Utilizing Eq. 4, the processing pipeline for a specified pixel at coordinates $(x,y)$ at any given time $t$ and $t^{\prime}$ can be delineated by Eq. 5.
-[…]
-Utilizing Eq.
-5, coupled with corresponding events, the intensity frame at a given time enables the computation of intensity frames for alternate times, facilitating video frame interpolation.
-However, events typically contain noise, and employing Eq.
-5 directly cannot produce optimal outcomes [45].
-Numerous studies [44, 62, 79] have demonstrated the effectiveness of employing neural networks to guide event-based modeling in robust frame interpolation.
-Additionally, the high temporal resolution of events aids in high spatial resolution conversion, a finding corroborated by prior research [38, 23].
-In conclusion, comprehending the event generation model reveals its substantial benefits for frame interpolation and VSR tasks, making it a natural guide for continuous space-time video super-resolution (C-STVSR) tasks.
-Appendix B More Details about Network Structure
-Owing to space constraints in the main text, this section provides a more detailed account of the network’s results.
-The description begins with the Regional Event Feature Extraction component, covering its preprocessing steps and the architecture of the Swin Transformer Encoder Blocks (STEB).
-Subsequently, the Holistic Event-Frame Feature Extraction is detailed, which, besides STEB, incorporates the Swin Transformer Decoder Block (STDB).
-In summary, this section elaborates on the preprocessing for both feature extraction components and details the structures of STEB and STDB.
-B.1 Regional Event Feature Extraction:
-This section delves into Regional Event Feature Extraction, denoted as $f_{re}$.
-$f_{re}$ receives Temporal Pyramid Representation (TPR) $E_{p}$ corresponding to timestamp $t$, formatted as $R^{L\times M_{p}\times H\times W}$.
-Here, $L$ signifies the number of TPR layers, while $M_{p}$ indicates the count of moments within each layer.
-$f_{re}$ involves two primary stages: preprocessing and feature extraction via STEB.
-Specifically, this preprocessing first increases its dimension to $C_{r}$ through a $1\times 1$ convolution operation.
-This preprocessing alters the TPR’s shape to $L\times C_{r}\times H\times W$.
-Feature extraction is then performed using STEB, where both the input and output retain their shape throughout the process.
-Swin Transformer Encoder Blocks (STEB): The STEB structure, proven effective in video super-resolution and frame interpolation [37, 35, 16], is adopted in our design.
-The input and output shapes of STEB are represented as $L\times C\times H\times W$ for clarity.
-Initially, for a window size of $M\times M$, specifically $4\times 4$ in our implementation, the input is partitioned into disjoint windows of $(M\times M)\times N\times(H/M)\times(W/M)\times C$ dimensions.
-Then, each window is compressed to form a feature map of shape $(M\times M)\times(N\times H\times W/M^{2})\times C$.
-Following this, Layer Normalization [2] and window-based multi-head self-attention [35, 16] are computed for each window, succeeded by further transformation via another Layer Normalization and a Multi-Layer Perception.
-Shifted window-based multi-head self-attention [35, 16] is then employed to establish cross-window connections.
-After applying one STEB structure, a second STEB is introduced with an identical configuration, except the input feature window is offset by $(M/2)\times(M/2)$.
-In total, four STEBs are utilized for comprehensive feature extraction in the Regional Event Feature Extraction.
-B.2 Holistic Event-Frame Feature Extraction:
-This section details the Holistic Event-Frame Feature Extraction ($f_{he}$) module.
-The $f_{he}$ module processes inputs $\boldsymbol{I_{in}}$ and $\boldsymbol{E_{v}}$ to generate the output $F^{g}$.
-Input $\boldsymbol{I_{in}}$ has a shape of $N\times 3\times H\times W$, whereas $\boldsymbol{E_{v}}$ is structured as $M\times H\times W$, with $M$ indicating the count of moments derived from the events.
-The output, $F^{g}$, has a structure of $C_{g}\times H\times W$.
-Initially, $f_{he}$ transforms $\boldsymbol{I_{in}}$ and $\boldsymbol{E_{v}}$.
-$\boldsymbol{I_{in}}$ and $\boldsymbol{E_{v}}$ are transformed into features with dimensions $N\times C\times H\times W$ and $N-1\times C\times H\times W$ using $1\times 1$ convolution and reshaping, followed by processing through a STEB module.
-The process then employs a combination of downsampling and STEB to extract features across multiple scales, each possessing a distinct view field.
-Notably, STEB maintains the same resolution and channel count in both its input and output.
-Next, the Swin Transformer Decoder (STDB) is utilized for fusing and decoding the dual-modal features.
-Each STDB block receives three types of inputs: the preceding STDB’s output and the outputs from the $\boldsymbol{I_{in}}$ and $\boldsymbol{E_{v}}$ encoder sections, all at a matching resolution.
-In practice, the outputs from the $\boldsymbol{I_{in}}$ and $\boldsymbol{E_{v}}$ encoders are concatenated, transformed into queries via an MLP layer, and then fused utilizing a multi-head attention mechanism.
-Notably, the first STDB module lacks a preceding STDB output, hence defaulting this input to zero.
-The final output, $F^{g}$, is derived from the last STDB’s output, post reshaping and $1\times 1$ convolution.
-In summary, $f_{he}$ employs a sequence of intricate yet efficacious transformation and fusion procedures, designed to extract multifaceted features from inputs $\boldsymbol{I_{in}}$ and $\boldsymbol{E_{v}}$ for subsequent tasks in video super-resolution and frame interpolation.
-Appendix C More Details about Experiments Setting
-C.1 More Details about Datasets
-To facilitate a comprehensive comparison between the frame-base [10, 8] and event-guided methods [38, 62, 61, 23], we employed four datasets: two simulated (Adobe240 [56] and GoPro [40]) and two real-world (HS-ERGB [61] and CED [51]) for our experiments.
-1) Adobe240 Dataset [56]: This dataset comprises 133 videos, each with a resolution of $720\times 1280$ and a frame rate of 240.
-We follow [10, 8, 70] to split this dataset into 100 training, 16 validation, and 17 test sets.
-Upon frame extraction, we employed the widely-used event simulation method vid2e [15], which accounts for real noise distribution, ensuring robust neural network training with enhanced generalization.
-In generating the inputs and ground truth of the network, we adopted and extended the previous works [8, 10] to accommodate multi-frame input.
-Specifically, input and ground truth (GT) frames are selected via sliding windows.
-We define the window size as $W$, the number of input frames as $N_{in}$, and the interval as $S$.
-They interrelate as: $W=(N_{in}-1)*(S+1)+1$ For instance, with 4 frames input at 7-frame intervals, 25 frames are chosen.
-The 1st, 9th, 17th, and 25th frames become the input after down-sampling, while frames 1-25 serve as the GT.
-We adopted two strategies in line with prior works [10, 8]: I) A fixed magnification set at 4×4$\times$4 × the input resolution, and II) A variable enlargement strategy, wherein the scaling factor is governed by a $\mathcal{U}(1,8)$ distribution.
-2) GoPro Dataset [40]: Both the GoPro and Adobe240 datasets share a resolution of $720\times 1280$ and a frame rate of 240 $fps$.
-However, the GoPro dataset is more compact, encompassing 22 training videos and 11 for testing.
-Owing to this, prior research [8, 10] predominantly utilized the GoPro test set for quantitative evaluations, neglecting its training set.
-In the interest of fairness, we adopted the same approach.
-3) BS-ERGB [61]: This dataset, captured with a spectroscope, comprises paired events and RGB frames from real-world scenarios.
-It has a resolution of $970\times 625$ with RGB frames captured at 28 fps.
-Of the 123 videos in the dataset, 47 are allocated for training, 19 for validation, and 26 for testing.
-Each video contains between 100 to 600 frames.
-Despite being a real-world dataset, its size is inadequate for standalone training of a frame interpolation network.
-In previous work, e.g.TimeLens++ [61], the Vime90k [72] simulation dataset was initially utilized for pre-training, followed by fine-tuning using this real-world dataset.
-We, on the other hand, opted for a model pre-trained on Adobe240 for our fine-tuning.
-4) CED dataset [51]: This is another real-world dataset in SR research where both frames and events exhibit a resolution of $346\times 260$.
-Following the previous research [23], we conducted preprocessing on this dataset.
-C.2 More Details about Implementation Details
-Our experiments encompass four datasets, compared against a variety of benchmark methods.
-To guarantee fairness in these comparisons, we provide additional details regarding our experimental approach.
-For the Adobe240 and GoPro datasets, our training and testing protocols adhere to methods established in prior research [8, 10].
-Specifically, our approach diverges from prior studies in handling the smaller-scale BS-ERGB dataset.
-Unlike previous methods that pre-trained on the larger Vime90k dataset before fine-tuning on BS-ERGB, we opted for a different strategy.
-Our method involves pre-training on the Adobe240 dataset.
-During this phase, we input four video frames, with a gap of seven frames between each pair.
-From a total of 25 frames, 20 are randomly chosen as the ground truth.
-In the pre-training phase, our focus is solely on modules related to frame insertion, omitting any upsampling procedures.
-Consequently, the spatial embedding aspect is excluded, and the decoding is performed using only a four-layer MLP decoder.
-For the super-resolution experiments on the CED dataset, we deviated from pre-training and instead trained our model exclusively on the CED dataset.
-A more detailed comparison of the 2×2$\times$2 × VSR is shown in Tab.
-8.
-This approach was chosen to maintain a fair comparison with other works [38, 23].
-Appendix D More Experiments Results and Analysis
-D.1 Stability of VFI in Various Timestamps:
-Figure 8: Timestamps and corresponding PSNR for each frame during 4×4$\times$4 × VSR and 7-skip VFI on the Adobe240 dataset [56].
-Figure 9: 4×4$\times$4 × VSR and 7-skip VFI visualization results. Please refer to the supplementary material for the video of this case.
-Our method demonstrates temporal stability, accurately estimating motion states at each time point during frame interpolation.
-Specifically, whether in the GoPro-Average and Adobe-Average of Tab.
-1, or the 12-skip and 16-skip tests of Tab.
-2, our method significantly outperforms previous methods by at least $1.2dB$ with 4×4$\times$4 × VSR.
+Moreover, even with two frames, our method also outperforms previous works (Chen et al, 2022, 2023b).
 Fig.
-8 shows the relationship between timestamps and PSNR, validating the greater stability of our method compared to VideoINR [10], especially around the $0.5s$ mark, where VideoINR experiences a notable decrease.
-This observation is also reflected in the visualization results of Fig.
-1 and Fig.
-5 in main paper and Fig.
-6 and Fig.
-9 in appendix.
-In Fig.
-6 and Fig.
-9, whether using real-world or simulated data, our method demonstrates superior temporal consistency, exhibiting fewer artifacts compared to previous methods such as VideoINR [10] and TimeLens [62].
-This showcases the real-world effectiveness of our method’s temporal stability.
-The main reason is that optical flow-based methods [10, 8, 62] suffer from instability in flow estimation at time far from the reference frames (at 00 and $1$ timestamp), which impacts the motion estimation.
-In contrast, this limitation does not affect our method and maintains higher stability throughout the temporal sequence.
-D.2 Video Super-resolution with 2×2$\times$2 × in CED dataset
-Table 8: Quantitative results (PSNR/SSIM) of the proposed our framework and other methods on the CED dataset for ×$\times 2$.
-Because the official training code is not available, * denoted values were acquired from the pre-trained model that the authors have released.
-Clip Name TDAN [60] SOF [64] RBPN[19] VideoINR [10]* E-VSR [23] EG-VSR [38] Ours
-people_dynamic_wave 35.83 / 0.9540 33.32 / 0.9360 40.07 / 0.9868 27.47 / 0.8229 41.08 / 0.9891 38.78 / 0.9794 41.50 / 0.9901
-indoors_foosball_2 32.12 / 0.9339 30.86 / 0.9253 34.15 / 0.9739 26.03 / 0.7766 34.77 / 0.9775 38.68 / 0.9750 42.17 / 0.9904
-simple_wires_2 31.57 / 0.9466 30.12 / 0.9326 33.83 / 0.9739 26.77 / 0.8321 34.44 / 0.9773 38.67 / 0.9815 42.21 / 0.9922
-people_dynamic_dancing 35.73 / 0.9566 32.93 / 0.9388 39.56 / 0.9869 27.36 / 0.8202 40.49 / 0.9891 39.06 / 0.9798 42.02 / 0.9913
-people_dynamic_jumping 35.42 / 0.9536 32.79 / 0.9347 39.44 / 0.9859 27.24 / 0.8183 40.32 / 0.9880 38.93 / 0.9792 42.09 / 0.9916
-simple_fruit_fast 37.75 / 0.9440 37.22 / 0.9390 40.33 / 0.9782 27.21 / 0.8456 40.80 / 0.9801 41.96 / 0.9821 43.96 / 0.9912
-outdoor_jumping_infrared_2 28.91 / 0.9062 26.67 / 0.8746 30.36 / 0.9648 26.88 / 0.8226 30.70 / 0.9698 38.03 / 0.9755 42.68 / 0.9902
-simple_carpet_fast 32.54 / 0.9006 31.83 / 0.8774 34.91 / 0.9502 24.21 / 0.5909 35.16 / 0.9536 36.14 / 0.9635 39.80 / 0.9853
-people_dynamic_armroll 35.55 / 0.9541 32.79 / 0.9345 40.05 / 0.9878 27.26 / 0.8193 41.00 / 0.9898 38.84 / 0.9787 41.99 / 0.9915
-indoors_kitchen_2 30.67 / 0.9323 29.61 / 0.9192 31.51 / 0.9551 26.44 / 0.7502 31.79 / 0.9586 37.68 / 0.9726 41.61 / 0.9901
-people_dynamic_sitting 35.09 / 0.9561 32.13 / 0.9367 39.03 / 0.9862 27.63 / 0.8230 39.97 / 0.9884 38.86 / 0.9810 41.99 / 0.9917
-average PSNR/SSIM 33.74 / 0.9398 31.84 / 0.9226 36.66 / 0.9754 26.77 / 0.7938 37.32 / 0.9783 38.69 / 0.9771 42.01 / 0.9905
-In Tab.
-8, our proposed method demonstrates the highest performance across various clips in the CED dataset, significantly outperforming previous methods, including EG-VSR.
-Our approach consistently achieves higher PSNR and SSIM values, indicating superior visual quality and reconstruction accuracy.
-For instance, on average, our method improves PSNR by 3.32 $dB$ and SSIM by 0.0134 compared to EG-VSR.
-These results highlight the effectiveness of our framework in handling complex scenes, showcasing its robustness and reliability in video super-resolution tasks.
-D.3 Holistic-Regional feature visualization
-Fig.
-10 (j, l, n) showcases TPR’s regional features, emphasizing its focus on local motion and highlighting its capability to capture local dynamics as a novel event representational form.
-Specifically, in a scenario where the camera is fixed while filming a basketball player, the background information is displayed in the Holistic Feature, whereas the motion information (such as the moving basketball) is emphasized in the Regional Feature.
-This indicates that our method can effectively capture different types of information through both Holistic and Regional features.
-For more cases, please refer to the example figure below and the videos in the supplementary materials.
-Figure 10: Feature visualization on real data [61]: (h) shows the holistic feature, $F^{g}$, derived from multiple frames and events; (j, l, n) depict the regional features ($F_{t}^{g}$), highlighting the capability to capture local motion.
-D.4 Time embedding feature visualization
-Figure 11: Visualization of Temporal Embedding.
-Figure (a) shows the original visualization of Temporal Embedding, while figure (b) displays the results after sorting.
-The sorting is based on the size of the first element, arranged in ascending order.
-We use MLP decoding, where the order is not crucial.
-However, to more clearly demonstrate the outcomes of Temporal Embedding learning, we have chosen to present the sorted results.
-Fig.11 illustrates the visualization of time embedding features.
-Compared to traditional sine-cosine embedding features, the learning-based approach performs better, as shown in Tab.
-5.
-The visualized learning-based embedding features not only demonstrate the capability to learn periodic positional representations but also provide a richer expression of exposure information.
-D.5 Bad case analysis
-We have observed that our method has certain limitations in some cases (Fig.
-9).
+14 compares the outputs of two models: one using two input frames (0, 1) and another using four input frames (-1, 0, 1, 2).
+With two input frames, the model struggles to capture complex motion, as seen in the less accurate predictions of the basketball’s trajectory in Fig.14 (b).
+In contrast, the four-frame model leverages additional temporal information to model longer-term motion, producing smoother and more accurate results, as shown in Fig.14 (c).
+The holistic feature visualizations further highlight the richer temporal dependencies captured by the four-frame model, allowing it to handle challenging motion scenarios more effectively.
+This analysis underscores the strength of our method in utilizing multi-frame inputs to improve temporal coherence and enhance interpolation quality, particularly for long-distance or rapid motion.
+Figure 15: Comparison of total and average time for $34\times$ frame interpolation by different methods.
+Our method takes less time than TimeLens (Tulyakov et al, 2021), but slightly more time than VideoINR (Chen et al, 2022).
+Inference time analysis In Fig.
+13, we analyze the inference times of three different methods.
+Both our method and VideoINR achieve an average frame time of less than 100 $ms$ for $34\times$ frame interpolation.
+In contrast, TimeLens (Tulyakov et al, 2021) has an average frame time of 187 $ms$, which is more than double that of our method.
+The tests were conducted on a high-performance computer, and each method was tested 30 times, with the final inference time being the average of these 30 trials.
+Bad case analysis We have observed that our method has certain limitations in some cases (Fig.
+8).
 For example, when restoring color information, although our model can accurately reconstruct the contours of objects, the color information is often distorted or missing.
@@
 We believe that with future advancements in color event technology, this problem will be effectively addressed.
-D.6 Inference time analysis
-In Fig.
-11, we analyze the inference times of three different methods.
-Both our method and VideoINR achieve an average frame time of less than 100 $ms$ for 34×34$\times$34 × frame interpolation.
-In contrast, TimeLens [62] has an average frame time of 187 $ms$, which is more than double that of our method.
-The tests were conducted on a high-performance computer, and each method was tested 30 times, with the final inference time being the average of these 30 trials.
-Figure 12: Comparison of total and average time for 34×34$\times$34 × frame interpolation by different methods.
-Our method takes less time than TimeLens [62], but slightly more time than VideoINR [10].
-Appendix E More Visualization Results
+5 Conclusion
+Our work introduced the first event-guided continuous space-time video super-resolution method. The main contributions are:
+  • Event temporal pyramid representation for capturing short-term dynamic motion;
+  • A feature extraction process combining holistic and regional features to manage motion dependencies;
+  • A spatiotemporal decoding process based on implicit neural representation, avoiding traditional optical flow and achieving stable frame interpolation through temporal-spatial embedding.
+Furthermore, the unique combination of event data with temporal pyramid representation allows our method to handle both fast-moving and stationary objects seamlessly, making it highly effective in real-world scenarios with complex motion.
+Our approach also outperforms previous methods in terms of stability, both in temporal and spatial resolution, as demonstrated by comprehensive experiments on multiple datasets.
+By integrating local and global motion features, our method offers enhanced adaptability, overcoming challenges such as ghosting artifacts and motion blur, commonly encountered in prior techniques.
+This not only sets a new benchmark in video super-resolution but also paves the way for more accurate and efficient event-based video processing in dynamic environments.
+In the future, we plan to extend this method to other fields, such as deblurring, to further enhance its applicability.
+Acknowledgements.
+This work was supported in part by the National Key R&D Program of China (Grant No.2023YFF0725001), in part by the National Natural Science Foundation of China (Grant No.92370204), in part by the guangdong Basic and Applied Basic Research Foundation (Grant No.2023B1515120057).
+in part by Guangzhou-HKUST(GZ) Joint Funding Program (Grant No.2023A03J0008), Education Bureau of Guangzhou Municipality.
+Supplementary Materials Preview:
 Additional videos have been included in the supplementary materials to provide a more comprehensive demonstration of our method’s visual results.
@@
   • 2-TimeLensPP-Ours-1: This video shows the performance of our method on real-world data sets and the visualization of features. Demonstrates that we effectively capture local motion.
-  • 3-Our-vs-Timelens: This video shows the results of comparing our method with Timelens [62].
-Figure 13: More visualization results on real-world data set [62].
-Figure 14: More visualization results on real-world data set [62].
-Figure 15: More visualization results on real-world data set [62].
-Figure 16: More visualization results on real-world data set [62].
-Figure 17: More visualization results on real-world data set [62].
-Figure 18: More visualization results on real-world data set [62].
-Figure 19: More visualization results on real-world data set [62].
-Figure 20: More visualization results on real-world data set [62].
+  • 3-Our-vs-Timelens: This video shows the results of comparing our method with Timelens.
+Figure 16: More visualization results on real-world data set (Tulyakov et al, 2021).
+Figure 17: More visualization results on real-world data set (Tulyakov et al, 2021).
+Figure 18: More visualization results on real-world data set (Tulyakov et al, 2021).
+Figure 19: More visualization results on real-world data set (Tulyakov et al, 2021).
+Figure 20: More visualization results on real-world data set (Tulyakov et al, 2021).
+Figure 21: More visualization results on real-world data set (Tulyakov et al, 2021).
+Figure 22: More visualization results on real-world data set (Tulyakov et al, 2021).